### v3

# Long time asymptotics for the KPII equation

ABSTRACT
  1. The long-time asymptotics of small Kadomtsev-Petviashvili II (KPII) solutions is derived using the inverse scattering theory and the stationary phase method.
1. Introduction
The Kadomtsev-Petviashvili II (KPII) equation
(1.1) $(-4u_{x_{3}}+u_{x_{1}x_{1}x_{1}}+6uu_{x_{1}})_{x_{1}}+3u_{{x_{2}}{x_{2}}}=0$
plays a significant role in plasma physics, water waves, and various other areas of mathematical physics. As one of the few physically relevant integrable systems in multiple spatial dimensions, the KPII equation has been the focus of extensive research. In particular, its global well-posedness and stability properties have been investigated through both partial differential equation (PDE) methods and the inverse scattering theory (IST). For a comprehensive overview of these developments, we refer the reader to the monograph by Klein and Saut [6].
Despite this progress, a complete description of the long-time behavior of KPII solutions remains largely open. Using PDE methods, the asymptotic behavior of small solutions to generalized KPII equations, excluding the KPII equation itself, has been investigated in works such as [4, 7]. In addition, the long-time asymptotics of the $x_{1}$-derivative of KPII solutions were studied in [3]. On the other hand, Kiselev formally derived the long-time $\color[rgb]{0,0,0}o(t^{-1})$ behavior of small KPII solutions using the IST [5]. However, his analysis relies on non-physical and non-generic assumptions, particularly the integrability of $(1+|\lambda|)s_{c}$ and boundedness of $\partial_{\lambda_{I}}s_{c},\,\partial^{2}_{\lambda_{I}}s_{c}$. Since the Lax operator associated with the KPII is the heat operator, the scattering data $s_{c}$ is naturally differentiable and decaying in $(\overline{\lambda}-\lambda,\overline{\lambda}^{2}-\lambda^{2})$, and the associated eigenfunction $m(x,\lambda)$ depends nontrivially on the entire complex plane. As a result, the assumptions imposed by Kiselev lead to highly degenerate scattering data along the real axis $\lambda_{I}=0$.
The goal of this paper is to rigorously establish the large-time asymptotic behavior of small solutions to the KPII equation, without imposing any non-physical assumptions. Our approach is based on IST [8], the representation formula (2.4) for the KPII solution $u$,
(1.2) $u(x)=$ $\ u_{1}(x)+u_{2,0}(x)+u_{2,1}(x),\quad x=(x_{1},x_{2},x_{3}),$
(1.3) $u_{1}(x)=$ $-\frac{1}{\pi i}\partial_{x_{1}}\iint e^{2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\ d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime},$
(1.4) $u_{2,0}(x)=$ $-\dfrac{1}{\pi i}\iint e^{2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})(\widetilde{m}(x,{\color[rgb]{0,0,0}\overline{\lambda}^{\prime}})-1)\ d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime},$
(1.5) $u_{2,1}(x)=$ $-\dfrac{1}{\pi i}\iint e^{2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\partial_{x_{1}}\widetilde{m}(x,{\color[rgb]{0,0,0}\overline{\lambda}^{\prime}})\ d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime},$
novel representation formulas for the Cauchy integrals (see Lemmas 4.2, 4.4, and 5.1), and the stationary phase method [2]. We eliminate non-physical conditions by performing integration by parts with respect to $\lambda^{\prime}_{I}$ or $\xi_{h}^{\prime\prime}$ in regimes where $|\lambda_{R}^{\prime}|<C$ or $|\lambda_{I}^{\prime}|>1/C$, and by carefully exploiting the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$ or $(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})$, which arise from taking the $x_{1}$-derivative in the representation formulas (1.4) or (1.5), in regimes where $|\lambda_{R}^{\prime}|>1/C$ or $|\lambda_{I}^{\prime}|<1/C$. See Section B for the definitions of $\widetilde{s}_{c},\,\widetilde{m},\,S_{0}$, $C$, $\lambda^{\prime}$, $\lambda_{R}^{\prime},\,\lambda_{I}^{\prime}$, and $\xi_{h}^{\prime\prime}$.
Our main result is as follows:
Theorem 1.
Let $a=\pm 3r^{2}=\frac{x_{2}^{2}-3x_{1}x_{3}}{3x_{3}^{2}}$, $r>0$, $t=-x_{3}$, ${|x_{2}|}/t<C$, and $\delta>0$ be any fixed constant. Suppose
(1.6) $\sum_{l_{1}+l_{2}\leq{\color[rgb]{0,0,0}8}}|\partial_{x_{1}}^{l_{1}}\partial_{x_{2}}^{l_{2}}(1+|x_{1}|+|x_{2}|)^{3}u_{0}(x_{1},x_{2})|_{L^{1}\cap L^{2}}<\infty,\quad|u_{0}(x_{1},x_{2})|_{L^{1}\cap L^{2}}\ll 1.$
Then, the solution $u$ to the Cauchy problem for (1.1) with initial data $u_{0}$ satisfies : as $t\to+\infty$,
[…]
    For $\color[rgb]{0,0,0}a>+\delta>0$,
    $u_{1}(x)\sim o(t^{-1}),\qquad u_{2,0}(x),\ u_{2,1}(x)\sim{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.$
[…]
    For $\color[rgb]{0,0,0}a<-\delta<0$,
    $u_{1}(x)\sim\frac{2ie^{i4\pi tr^{3}}}{3t}s_{c}(-\frac{x_{2}}{3x_{3}}+ir)-\frac{2ie^{-i4\pi tr^{3}}}{3t}s_{c}(-\frac{x_{2}}{3x_{3}}-ir)+o(t^{-1}),$
    $u_{2,0}(x),\ u_{2,1}(x)\sim o(t^{-8/9}).$
Here, $s_{c}(\lambda)$ denotes the scattering data of $u_{0}$, $a$ characterizes the stationary points in the phase, and $t$ corresponds to the direction of KPII propagation.
[…]
The paper is organized as follows. In Section 2, we present preliminary materials, including the IST for the KPII equation and an introduction to the stationary phase method.
In Section 3, we first establish the $\lambda^{\prime}$-derivative estimates for the scattering data, which, together with Theorem 1, form the basis of the asymptotic analysis. We then derive the asymptotic behavior of $u_{1}$ by applying the stationary phase method near the stationary points and using integration by parts away from them.
[…]
In Subsection 4.2 and 4.3, we determine asymptotic behavior of $u_{2,0}$ in the regimes $a\gtrless 0$, respectively. This is achieved by refining the decomposition of the representation formulas, establishing the integrability of $\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}$ or $(1+|\lambda^{\prime}|)\widetilde{s}_{c}$ in various regimes, discarding terms with rapidly decaying amplitudes, and using several key tools: smallness of the integration domains, the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, integration by parts, and the estimates developed in Subsection 4.1.
In Section 5, we adapt the approach from Section 4 to investigate the Cauchy integrals $\partial_{x_{1}}\widetilde{\left(\mathcal{C}T\right)^{n}1}$ and derive the asymptotic behavior of $u_{2,1}$. To facilitate integration by parts without imposing additional conditions on $\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}$ and $(1+|\lambda^{\prime}|)\widetilde{s}_{c}$, particular care is needed, and the argument becomes more involved.
In Appendices A and B, we provide a key estimate used in the derivation of the new representation formulas, along with a list of symbols used throughout the paper.
Acknowledgments. I am grateful to J.-C. Saut for suggesting the asymptotic problem of the KP equations. I would like to express my special thanks to Jiaqi Liu for the inspiring discussions that led to the discovery of novel representation formulas for the Cauchy integrals. This research was supported by NSC 113-2115-M-001-007-.
2. Preliminaries
2.1. The IST for KPII equations
Denote $x=(x_{1},x_{2},x_{3})$, $l=(l_{1},l_{2},l_{3})$, $\partial_{x}^{l}=\partial_{x_{1}}^{l_{1}}\partial_{x_{2}}^{l_{2}}\partial_{x_{3}}^{l_{3}}$, $|l|=|l_{1}|+|l_{2}|+|l_{3}|$, $\widehat{f}(\xi)=\widehat{f}(\xi_{1},\xi_{2})=\iint f(x)e^{-2\pi i(x_{1}\xi_{1}+x_{2}\xi_{2})}dx_{1}dx_{2}$, $C$ a uniform constant that is independent of $x$, $\lambda$, and $\mathfrak{M}^{p,q}=\{f{\color[rgb]{0,0,0}(x_{1},x_{2})}:\sum_{|l|\leq{q}}|{\color[rgb]{0,0,0}\partial_{x_{1}}^{l_{1}}\partial_{x_{2}}^{l_{2}}}(1+{\color[rgb]{0,0,0}|x_{1}|+|x_{2}|})^{p}f|_{L^{1}\cap L^{2}}<\infty\}$.
By establishing an IST, Wickerhauser solved the Cauchy problem of the KPII equation with a vacuum background:
Theorem 2 (The Cauchy Problem [8]).
Let $q\geq{\color[rgb]{0,0,0}8}$. If the initial data $u_{0}\in\mathfrak{M}^{0,q}$ satisfies
(2.1) $u_{0}(x_{1},x_{2})=\overline{u_{0}(x_{1},x_{2})},\qquad|u_{0}|_{\mathfrak{M}^{0,0}}\ll 1.$
Then, we can construct the forward scattering transform:
(2.2) $\begin{split}\mathcal{S}:u_{0}\mapsto s_{c}(\lambda)=&\frac{\operatorname{sgn}(\lambda_{I})}{2\pi i}\left[{u_{0}(\cdot)m_{0}(\cdot,\lambda)}\right]^{\wedge}(\frac{\overline{\lambda}-\lambda}{2\pi i},\frac{\overline{\lambda}^{2}-\lambda^{2}}{2\pi i})\equiv\frac{\operatorname{sgn}(\lambda_{I})}{2\pi i}\left[{u_{0}m_{0}}\right]^{\wedge}(\xi_{1},\xi_{2}),\end{split}$
satisfying the algebraic and analytic constraints:
(2.3) $s_{c}(\lambda)=\overline{s_{c}(\overline{\lambda})},\quad|(1+|\xi|)^{q}s_{c}(\lambda(\xi))|_{L^{\infty}\cap L^{2}(d\xi_{1}d\xi_{2})}\leq{C|u_{0}|_{\mathfrak{M}^{0,q}}}.$
Here $m_{0}$ solves the boundary value problem of the Lax equation:
(2.4) $(-\partial_{x_{2}}+\partial_{x_{1}}^{2}+2\lambda\partial_{x_{1}}+u_{0}(x_{1},x_{2}))m_{0}(x_{1},x_{2},\lambda)=0,\ \ \ \lim_{|x|\to\infty}m_{0}(x_{1},x_{2},\lambda)=1,$
Moreover, the solution $u$ to the KPII Cauchy problem is given by:
(2.5) $u(x)=-\frac{1}{\pi i}\partial_{x_{1}}\iint Tm\ d\overline{\zeta}\wedge d\zeta,$
satisfying
(2.6) $u(x)=\overline{u(x)},\quad|u|_{\mathfrak{M}^{0,q-{\color[rgb]{0,0,0}4}}}\leq C|u_{0}|_{\mathfrak{M}^{0,q}}.$
Here $m$ solves the Cauchy integral equation:
(2.7) ${m}(x,\lambda)=1+\mathcal{C}Tm(x,\lambda),\quad m_{0}(x_{1},x_{2},\lambda)=m(x_{1},x_{2},0,\lambda)$
with $\mathcal{C}$ being the Cauchy integral operator, and $T$ the continuous scattering operator:
(2.8) $\mathcal{C}\phi(x,\lambda)\equiv$ $-\frac{1}{2\pi i}\iint_{{\mathbb{C}}}\frac{\phi(x,\zeta)}{\zeta-\lambda}d\overline{\zeta}\wedge d\zeta,$
(2.9) $T\phi(x,\lambda)\equiv$ $e^{(\overline{\lambda}-\lambda)x_{1}+(\overline{\lambda}^{2}-\lambda^{2})x_{2}+(\overline{\lambda}^{3}-\lambda^{3})x_{3}}{s}_{c}(\lambda)\phi(x,\overline{\lambda}).$
2.2. The stationary points
Building upon Theorem 2, we are going to investigate the long-time asymptotic behavior of the KPII solution using the stationary phase method (cf [2] for the corresponding analysis in the KPI case). The natural coordinates for applying this method are the variables $(\zeta_{R}^{\prime},\zeta_{I}^{\prime})$ introduced in (2.12). To motivate their use, we define :
(2.10) $\begin{split}t_{1}=\frac{x_{1}}{t},\quad t_{2}=\frac{x_{2}}{t},&\quad t=-x_{3},\\ 2\pi i\xi_{1}=\overline{\zeta}-\zeta,&\quad 2\pi i\xi_{2}=\overline{\zeta}^{2}-\zeta^{2},\\ \zeta=\frac{\xi_{2}}{2\xi_{1}}-i\pi\xi_{1}=\zeta_{R}+i\zeta_{I},&\quad d\overline{\zeta}\wedge d\zeta=2i\,d\zeta_{R}d\zeta_{I}=\frac{i\pi}{|\xi_{1}|}d\xi_{1}d\xi_{2}.\end{split}$
and the phase function $\mathbb{S}_{0}$ by
(2.11) $\mathbb{S}_{0}(t_{1},t_{2};\zeta(\xi))\equiv\frac{(\bar{\zeta}-\zeta)x_{1}+(\bar{\zeta}^{2}-\zeta^{2})x_{2}+(\bar{\zeta}^{3}-\zeta^{3})x_{3}}{2\pi it}.$
Notice that due to the propagation of the KPII equation (1.1), we will investigate the asymptotic of the KPII solution $u(x)$ as $t\to\infty$.
To simplify the computation by eliminating quadratic terms, we introduce :
(2.12) $(\zeta,\overline{\zeta})=(\zeta^{\prime}+\frac{t_{2}}{3},\overline{\zeta}^{\prime}+\frac{t_{2}}{3}),$ $\quad(\xi^{\prime}_{1},\xi^{\prime}_{2})=(\xi_{1},\xi_{2}-\frac{2t_{2}}{3}\xi_{1}),$
$2\pi i\xi^{\prime}_{1}=\overline{\zeta}^{\prime}-\zeta^{\prime},$ $\quad 2\pi i\xi_{2}^{\prime}=\overline{\zeta}^{\prime 2}-\zeta^{\prime 2},$
$\zeta^{\prime}=\frac{\xi_{2}^{\prime}}{2\xi_{1}^{\prime}}-i\pi\xi_{1}^{\prime}=\zeta^{\prime}_{R}+i\zeta^{\prime}_{I},$ $\quad d\overline{\zeta}^{\prime}\wedge d\zeta^{\prime}=2i\,d\zeta_{R}^{\prime}d\zeta_{I}^{\prime}=\frac{i\pi}{|\xi_{1}^{\prime}|}d\xi^{\prime}_{1}d\xi_{2}^{\prime},$
$\partial_{\zeta^{\prime}_{I}}=-\frac{1}{\pi}\partial_{\xi^{\prime}_{1}}-\frac{1}{\pi}\frac{\xi^{\prime}_{2}}{\xi^{\prime}_{1}}\partial_{\xi^{\prime}_{2}},$ $\quad\partial_{\zeta^{\prime}_{R}}=2\xi^{\prime}_{1}\partial_{\xi^{\prime}_{2}},$
which induces the definition, estimates
(2.13) $\begin{gathered}f(\zeta)=f(\zeta^{\prime}+\frac{t_{2}}{3})\equiv\widetilde{f}(\zeta^{\prime}),\\ {\color[rgb]{0,0,0}(|\zeta_{R}|^{l_{1}}+|\zeta_{I}|^{l_{2}})|\partial_{\zeta_{R}}^{j_{1}}\partial_{\zeta_{I}}^{j_{2}}s_{c}|\sim(|\zeta_{R}^{\prime}|^{l_{1}}+|\zeta_{I}^{\prime}|^{l_{2}})|\partial_{\zeta^{\prime}_{R}}^{j_{1}}\partial_{\zeta^{\prime}_{I}}^{j_{2}}\widetilde{s}_{c}|},\quad\zeta^{\prime}_{I}\neq 0,\end{gathered}$
by $|t_{2}|<C$, and changes the phase function to
(2.14) $\begin{split}\mathbb{S}_{0}(t_{1},t_{2};\zeta(\xi))\equiv&\frac{1}{2\pi i}[a(\overline{\zeta}^{\prime}-\zeta^{\prime})-(\overline{\zeta}^{\prime 3}-\zeta^{\prime 3})]=-\frac{1}{\pi}(a\zeta_{I}^{\prime}+\zeta_{I}^{\prime 3}-3\zeta_{I}^{\prime}\zeta_{R}^{\prime 2})\\ =&a\xi_{1}^{\prime}+\pi^{2}\xi_{1}^{\prime 3}-\frac{3}{4}\frac{\xi_{2}^{\prime 2}}{\xi_{1}^{\prime}}\equiv S_{0}(a;\zeta^{\prime}(\xi^{\prime})),\end{split}$
with
(2.15) $a=t_{1}+\frac{1}{3}t_{2}^{2}.$
Definition 1.
Let the phase function $S_{0}(a;\zeta^{\prime})$ be defined by (2.14) and (2.15). Define
  • For $a<0$, the stationary points of $S_{0}$ are purely imaginary:
    (2.16) $\zeta^{\prime}_{R}=0,\quad\zeta^{\prime}_{I}=\pm\sqrt{\frac{-a}{3}}\equiv\pm r,\quad r>0.$
  • For $a>0$, the stationary points of $S_{0}$ are purely real:
    (2.17) $\zeta^{\prime}_{R}=\pm\sqrt{\frac{a}{3}}\equiv\pm r,\quad\zeta^{\prime}_{I}=0,\quad r>0.$
3. Long time asymptotics of $u_{1}(x)$
3.1. Estimates on scattering data
In this subsection, we provide estimates of derivatives of the scattering data.
Notice that, for fixed $\lambda$ [8, Equation (II.9)],
(3.1) $m_{0}(x_{1},x_{2},\lambda)\equiv$ $1+G_{u_{0}}m_{0},$
where
(3.2) $\begin{gathered}G_{u_{0}}f=1+\iint e^{2\pi i(x_{1}\xi_{1}+x_{2}\xi_{2})}\frac{\left[u_{0}f\right]^{\wedge}(\xi_{1},\xi_{2},\lambda)}{p_{\lambda}(\xi_{1},\xi_{2})}d\xi_{1}d\xi_{2},\\ p_{\lambda}(\xi_{1},\xi_{2})=(2\pi i\xi_{1}+\lambda)^{2}-(2\pi i\xi_{2}+{\lambda}^{2}).\end{gathered}$
Lemma 3.1.
(3.3) $|G_{u_{0}}f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}|f|_{L^{\infty}},$
(3.4) $|\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}},\ |\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}},$
and, for $j\geq 2$,
(3.5) $|\left[\partial^{j}_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|\lambda_{I}|)^{j-1}(1+|x_{2}|)^{j-1}|u_{0}|_{\mathfrak{M}^{j-1,0}}|f|_{L^{\infty}}+C|\partial^{j-1}_{\lambda_{R}}\left(\lambda_{I}G_{u_{0}}f\right)|,$
(3.6) $|\left[\partial^{j}_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|\lambda|)^{j-1}(1+|x_{2}|)^{j-1}|u_{0}|_{\mathfrak{M}^{j-1,0}}|f|_{L^{\infty}}+C|\partial^{j-1}_{\lambda_{I}}\left(\lambda_{I}G_{u_{0}}f\right)|,$
(3.7) $|\left[\partial_{\lambda_{R}}\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|{\color[rgb]{0,0,0}\lambda_{R}}|)(1+|x_{2}|)|u_{0}|_{\mathfrak{M}^{1,0}}|f|_{L^{\infty}}+C|\partial_{\lambda_{I}}\left(\lambda_{I}G_{u_{0}}f\right)|.$
Proof.
We can prove (3.3) by applying the estimates [8]:
(3.8) $\begin{gathered}\left|\frac{1}{p_{\lambda}}\right|_{L^{1}(\Omega_{\lambda},d\xi_{1}d\xi_{2})}\leq\frac{C}{(1+|\lambda_{I}|^{2})^{1/2}},\quad\left|\frac{1}{p_{\lambda}}\right|_{L^{2}(\Omega_{\lambda}^{c},d\xi_{1}d\xi_{2})}\leq\frac{C}{(1+|\lambda_{I}|^{2})^{1/4}},\end{gathered}$
where $\Omega_{\lambda}=\{(\xi_{1},\xi_{2})\in{\mathbb{R}}^{2}:|p_{\lambda}(\xi_{1},\xi_{2})|<1\}$.
For $(x_{2}-x_{2}^{\prime})\neq 0$, via the Fourier inversion formula, the convolution, and the residue theorem, the operator $G$ can be written as [1]:
(3.9) $G_{u_{0}}f(x_{1},x_{2},\lambda)$
$=$ $\iint\left[\frac{1}{p_{\lambda}(\xi_{1},\xi_{2})}\right]^{\vee_{\xi_{1},\xi_{2}}}(x_{1}-x^{\prime}_{1},x_{2}-x^{\prime}_{2})\left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})dx^{\prime}_{1}dx^{\prime}_{2}$
$=$ $\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\iint\frac{e^{2\pi i((x_{1}-x_{1}^{\prime})\xi_{1}+(x_{2}-x_{2}^{\prime})\xi_{2})}}{p_{\lambda}(\xi_{1},\xi_{2})}d\xi_{1}d\xi_{2}$
$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\iint\frac{e^{2\pi i((x_{1}-x_{1}^{\prime})\xi_{1}+(x_{2}-x_{2}^{\prime})\xi_{2})}}{\xi_{2}-(2\pi i\xi_{1}^{2}+2\xi_{1}\lambda)}d\xi_{1}d\xi_{2}$
$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})},$
where $\theta(s)$ is the Heaviside function. Hence,
(3.10) $\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}\left(4\pi i(x_{2}-x_{2}^{\prime})\xi_{1}\right)$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}{\color[rgb]{0,0,0}+}\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi\lambda_{I})G_{u_{0}}f.$
Along with (3.3), yields
$|\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}}.$
Similarly,
(3.11) $\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
$\times$ $\left[\partial_{\lambda_{I}}\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\right]e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
- $\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\cdot[-4\pi\xi_{1}(x_{2}-x_{2}^{\prime})]e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})},$
and
$|\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}}.$
Hence the proof of (3.4) is completed.
Moreover, from (3.10),
$\left[\partial_{\lambda_{R}}^{2}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}\partial^{2}_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
- $\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}(4\pi\lambda_{I})\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
${\color[rgb]{0,0,0}+}$ $\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi i\lambda_{I})\partial_{\lambda_{R}}\left(Gf\right).$
Hence the proof of (3.5)-(3.7) can be justified by bootstrapping (3.10) or (3.11).
Lemma 3.2.
Suppose that (2.1) holds.
(3.12) $|m_{0}-1|\leq C|u_{0}|_{\mathfrak{M}^{0,0}},\ \ |\partial_{\lambda_{R}}m_{0}|,\ |\partial_{\lambda_{I}}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|),$
(3.13) $|\partial_{\lambda_{R}}^{j}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{j-1,0}}(1+|x_{2}|)^{j-1}(1+|\lambda_{I}|)^{j},\ j\leq 3,$
(3.14) $|\partial_{\lambda_{I}}^{l}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{1-l,0}}(1+|x_{2}|)^{l-1}(1+|\lambda|)^{l},\ 2\leq l\leq 3,$
(3.15) $|\partial_{\lambda_{R}}\partial_{\lambda_{I}}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{1,0}}(1+|x_{2}|)(1+|{\color[rgb]{0,0,0}\lambda}|)^{2}.$
Proof.
Estimate (3.12) follows from Lemma 3.1 and
(3.16) $\begin{gathered}m_{0}=(1-G_{u_{0}})^{-1}1,\\ \partial_{\lambda_{R}}m_{0}=(1-G_{u_{0}})^{-1}\left[\partial_{\lambda_{R}}G_{u_{0}}\right]m_{0},\quad\partial_{\lambda_{I}}m_{0}=(1-G_{u_{0}})^{-1}\left[\partial_{\lambda_{I}}G_{u_{0}}\right]m_{0}.\end{gathered}$
Applying the bootstrapping argument to (3.16) and Lemma 3.1, we can justify (3.13)-(3.15). ∎
Combining Lemma 3.2 with (2.2), (2.10), and (2.12), we further deduce
Lemma 3.3.
Suppose the assumption of Theorem 1 holds for $\mathfrak{M}^{2,q}$.
(3.17) $|\partial_{\lambda^{\prime}_{R}}^{j}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{j-1,0}}(1+|\lambda^{\prime}_{I}|)^{j},\ 1\leq j\leq 3,$
(3.18) $|\partial_{\lambda^{\prime}_{I}}^{l}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{l-1,0}}(1+|\lambda^{\prime}|)^{l},\ \ 1\leq l\leq 3,$
(3.19) $|\partial_{\lambda^{\prime}_{R}}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{1,0}}(1+|\lambda^{\prime}|)^{2}.$
3.2. Long time asymptotics of $u_{1}(x)$
Throughout this subsection, $a$, $r$, $_{i}, tt$ are as defined in Definition 1 and the assumption of Theorem 1 holds. Let $\psi$ be a non negative smooth cutoff function such that $\psi(s)=1$ for $|s|\leq\frac{1}{2}$ and $\psi(s)=0$ for $|s|\geq 1$. Given $w_{0}\in\mathbb{R}$, define
(3.20) $\psi_{r,w_{0}}(s)=\psi\left(\frac{16(s-w_{0})}{r}\right)+\psi\left(\frac{16(s+w_{0})}{r}\right).$
Let
(3.21) $\chi(\zeta^{\prime})=\left\{{\begin{array}[]{ll}\psi_{r,r}(\zeta^{\prime}_{R})\psi_{r,0}(\zeta^{\prime}_{I}),&\textit{ for $a>0$,}\\ \psi_{r,r}(\zeta^{\prime}_{I})\psi_{r,0}(\zeta^{\prime}_{R}),&\textit{ for $a<0$.}\end{array}}\right$.
Decompose the linearized term $u_{1}(x)$, defined by (1.3), into
(3.22) $\qquad u_{1}(x)=u_{1,1}(x)+u_{1,2}(x),$
(3.23) $u_{1,1}(x)=$ $-\dfrac{1}{\pi i}\iint\widetilde{s}_{c}(\zeta^{\prime})e^{2\pi itS_{0}}(\overline{\zeta}^{\prime}-\zeta^{\prime})\chi(\zeta^{\prime})\ d\overline{\zeta}^{\prime}\wedge d\zeta^{\prime},$
(3.24) $u_{1,2}(x)=$ $-\dfrac{1}{\pi i}\iint\widetilde{s}_{c}(\zeta^{\prime})e^{2\pi itS_{0}}(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi(\zeta^{\prime}))\ d\overline{\zeta}^{\prime}\wedge d\zeta^{\prime}.$
The integration by parts approach is based on the following key estimate on the phase function:
Lemma 3.4.
On the support of $1-\chi(\zeta^{\prime})$, the phase function $S_{0}$ satisfies:
(3.25) $|\nabla S_{0}|\equiv|(\partial_{\zeta^{\prime}_{R}}S_{0},\partial_{\zeta^{\prime}_{I}}S_{0})|\geq$ $C(|a|+|\zeta^{\prime}|^{2}),$
(3.26) $|\Delta S_{0}|\equiv|{\color[rgb]{0,0,0}(\partial_{\zeta^{\prime}_{R}}^{2}S_{0},\partial_{\zeta^{\prime}_{I}}^{2}S_{0})}|\leq$ $C|\zeta^{\prime}|.$
Proof.
We have
(3.27) $\begin{split}\partial_{\zeta^{\prime}_{R}}S_{0}=&+\frac{6}{\pi}\zeta^{\prime}_{R}\zeta^{\prime}_{I},\quad\partial_{\zeta^{\prime}_{I}}S_{0}=+\frac{1}{\pi}(-a+3({\zeta_{R}^{\prime}}^{2}-{\zeta_{I}^{\prime}}^{2})).\end{split}$
Therefore (3.26) is justified and
(3.28) $|\partial_{\zeta^{\prime}_{R}}S_{0}|^{2}+|\partial_{\zeta^{\prime}_{I}}S_{0}|^{2}\geq$ ${\zeta^{\prime}_{R}}^{4}+2{\zeta^{\prime}_{R}}^{2}{\zeta^{\prime}_{I}}^{2}+({\zeta^{\prime}_{I}}^{2}+\frac{a}{3})^{2},\qquad a<0,$
(3.29) $|\partial_{\zeta^{\prime}_{R}}S_{0}|^{2}+|\partial_{\zeta^{\prime}_{I}}S_{0}|^{2}\geq$ ${\zeta^{\prime}_{I}}^{4}+2{\zeta^{\prime}_{R}}^{2}{\zeta^{\prime}_{I}}^{2}+({\zeta^{\prime}_{R}}^{2}-\frac{a}{3})^{2},\qquad a>0.$
Since proofs are identical. We only give the proof of (3.25) for $a<0$ for simplicity. By assumption (1), if $\psi_{r,r}(\zeta^{\prime}_{I})=1$, then $\psi_{r,0}(\zeta^{\prime}_{R})\neq 1$. Namely,
(3.30) $\frac{||\zeta^{\prime}_{I}|-r|}{r}\leq\frac{1}{32}<\frac{|\zeta^{\prime}_{R}|}{r},$
along with $r\sim\pm\sqrt{\frac{-a}{3}}$ and (3.28), implies that
(3.31) $|\partial_{\zeta^{\prime}_{R}}S_{0}|^{2}+|\partial_{\zeta^{\prime}_{I}}S_{0}|^{2}\geq C({\zeta^{\prime}_{R}}^{4}+{\zeta^{\prime}_{I}}^{4})\geq C({\zeta^{\prime}_{R}}^{4}+{\zeta^{\prime}_{I}}^{4}+a^{2}).$
On the other hand, if $\psi_{r,r}(\zeta^{\prime}_{I})\neq 1$, then there exists $C>1$ such that
(3.32)  $\textit{ either $|\zeta^{\prime}_{I}|\leq\frac{1}{C}r$ or $|\zeta^{\prime}_{I}|\geq Cr$ holds}.$
Applying (3.28), we have
(3.33) $\begin{split}|\partial_{\zeta^{\prime}_{R}}S_{0}|^{2}+|\partial_{\zeta^{\prime}_{I}}S_{0}|^{2}\geq&C({\zeta^{\prime}_{R}}^{4}+a^{2})\geq C({\zeta^{\prime}_{R}}^{4}+{\zeta^{\prime}_{I}}^{4}+a^{2}),\quad\,|\zeta^{\prime}_{I}|\leq\frac{1}{C}r,\\ |\partial_{\zeta^{\prime}_{R}}S_{0}|^{2}+|\partial_{\zeta^{\prime}_{I}}S_{0}|^{2}\geq&C({\zeta^{\prime}_{R}}^{4}+{\zeta^{\prime}_{I}}^{4})\geq C({\zeta^{\prime}_{R}}^{4}+{\zeta^{\prime}_{I}}^{4}+a^{2}),\quad|\zeta^{\prime}_{I}|\geq Cr.\end{split}$
Proposition 3.1.
Suppose the assumption of Theorem 1 holds and $|a|>\delta$. Then
(3.34) $|u_{1,2}(x)|\sim o(t^{-1}).$
Proof.
Integration by parts, applying Lemmas 3.3, 3.4, we have
(3.35) $|u_{1,2}(x)|\leq$ $\frac{C}{t}|\iint e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3}-3\zeta^{\prime}_{I}{\zeta^{\prime}_{R}}^{2})}\nabla\cdot\left(\widetilde{s}_{c}(\zeta^{\prime})(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)d\zeta^{\prime}_{R}d\zeta^{\prime}_{I}|,$
with
(3.36) $|\nabla\cdot\left(\widetilde{s}_{c}(\zeta^{\prime})(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)|_{L^{1}(d\zeta^{\prime}_{R}d\zeta^{\prime}_{I})}<C.$
Here note that discontinuity of $\tilde{s}_{c}$ at $\zeta_{I}^{\prime}=0$ can be neglected thanks to the factor $(\overline{\zeta}^{\prime}-\zeta^{\prime})$.
Setting $\tilde{\zeta}_{R}=\zeta^{\prime}_{I}{\zeta^{\prime}_{R}}^{2}$, for $\zeta^{\prime}_{R}\gtrless 0$, $\zeta_{I}\gtrless 0$,
(3.37) $|u_{1,2}(x)|$
$\leq$ $\frac{C}{t}|\int_{0}^{\infty}\int_{0}^{\infty}e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3}-3\tilde{\zeta}_{R})}\nabla\cdot\left(\widetilde{s}_{c}(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)\frac{\partial(\zeta^{\prime}_{R},\zeta^{\prime}_{I})}{\partial(\tilde{\zeta}_{R},\zeta^{\prime}_{I})}d\tilde{\zeta}_{R}d\zeta^{\prime}_{I}|$
$+$ $\frac{C}{t}|\int_{-\infty}^{0}\int_{-\infty}^{0}e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3}-3\tilde{\zeta}_{R})}\nabla\cdot\left(\widetilde{s}_{c}(\zeta)(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)\frac{\partial(\zeta^{\prime}_{R},\zeta^{\prime}_{I})}{\partial(\tilde{\zeta}_{R},\zeta^{\prime}_{I})}d\tilde{\zeta}_{R}d\zeta^{\prime}_{I}|$
where
(3.38) $\qquad|\nabla\cdot\left(\widetilde{s}_{c}(\zeta^{\prime})(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)\times\frac{\partial(\zeta^{\prime}_{R},\zeta^{\prime}_{I})}{\partial(\tilde{\zeta}_{R},\zeta^{\prime}_{I})}|_{L^{1}(d\tilde{\zeta}_{R}d\zeta^{\prime}_{I})}\\ =|\nabla\cdot\left(\widetilde{s}_{c}(\zeta^{\prime})(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)|_{L^{1}(d\zeta^{\prime}_{R}d\zeta^{\prime}_{I})}<C.\qquad$
Therefore (3.34) follows from Fubini’s theorem and the Riemann-Lebesgue lemma.
Proposition 3.2.
Suppose the assumption of Theorem 1 holds for $\mathfrak{M}^{2,q}$. Then, as $t\to+\infty$:
(3.39) $u_{1,1}(x)\sim\frac{2ie^{i4\pi tr^{3}}}{3t}\widetilde{s}_{c}(+ir)-\frac{2ie^{-i4\pi tr^{3}}}{3t}\widetilde{s}_{c}(-ir)+\mathcal{O}(t^{-4/3}),\ \textit{for $\color[rgb]{0,0,0}a<-\delta<0$},$
(3.40) $u_{1,1}(x)\sim\mathcal{O}(t^{-4/3}),\qquad\textit{for $\color[rgb]{0,0,0}a>+\delta>0$}.$
Proof.
[…]
(3.41) $u_{1,1}(x)=$ $-\frac{2}{\pi}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}\psi_{r,r}(\zeta^{\prime}_{I})(\overline{\zeta}^{\prime}-\zeta^{\prime})\int d\zeta^{\prime}_{R}\ \ e^{-\pi it(-\frac{6}{\pi}\zeta^{\prime}_{I}){\zeta^{\prime}_{R}}^{2}}\psi_{r,0}(\zeta^{\prime}_{R})\widetilde{s}_{c}(\zeta^{\prime}).$
Define the Fourier transforms as $\widehat{\phi}(\eta^{\prime}_{R},\eta^{\prime}_{I})=\phi^{\wedge_{\zeta^{\prime}_{R}}}\phi^{\wedge_{\zeta^{\prime}_{I}}}$ where
(3.42) $\begin{gathered}\phi^{\wedge_{\zeta^{\prime}_{R}}}(\eta^{\prime}_{R},\zeta^{\prime}_{I})=\int e^{-2\pi i\zeta^{\prime}_{R}\eta^{\prime}_{R}}\phi(\zeta^{\prime}_{R},\zeta^{\prime}_{I})d\zeta^{\prime}_{R},\\ \phi^{\wedge_{\zeta^{\prime}_{I}}}(\zeta^{\prime}_{R},\eta^{\prime}_{I})=\int e^{-2\pi i\zeta^{\prime}_{I}\eta^{\prime}_{I}}\phi(\zeta^{\prime}_{R},\zeta^{\prime}_{I})d\zeta^{\prime}_{I}.\end{gathered}$
[…]
$|\partial^{j}_{\zeta^{\prime}_{R}}f|_{L^{2}(d\zeta^{\prime}_{R})}<C,\quad|(1+|\eta^{\prime}_{R}|^{3})f^{\wedge_{\zeta^{\prime}_{R}}}|_{L^{2}(d\eta^{\prime}_{R})}<C,\quad|(1+{\eta^{\prime}_{R}}^{2})f^{\wedge_{\zeta^{\prime}_{R}}}|_{L^{1}(d\eta^{\prime}_{R})}<C.$
Hence we can apply the stationary phase theorem to get
(3.43) $u_{1,1}=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}e^{\pi i\frac{\operatorname{sgn}(\zeta^{\prime}_{I})}{4}}\frac{1}{\sqrt{|\frac{6}{\pi}\zeta^{\prime}_{I}|}}\int d\eta^{\prime}_{R}\left(1+\mathcal{O}(\frac{{\eta^{\prime}}^{2}_{R}}{t|\zeta^{\prime}_{I}|})\right)f^{\wedge_{\zeta^{\prime}_{R}}}(\eta^{\prime}_{R},\zeta^{\prime}_{I})$
$=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}-{\zeta^{\prime}_{I}}^{3})}e^{+\pi i\frac{\operatorname{sgn}(\zeta^{\prime}_{I})}{4}}\frac{1}{\sqrt{|\frac{6}{\pi}\zeta^{\prime}_{I}|}}\psi_{r,r}(\zeta^{\prime}_{I})(\overline{\zeta}^{\prime}-\zeta^{\prime})\widetilde{s}_{c}(0,\zeta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}}).$
[…]
(3.44) $|\partial^{j}_{\zeta^{\prime}_{I}}g|_{L^{2}(d\zeta^{\prime}_{I})}<C,\quad|(1+{\eta^{\prime}_{I}}^{3})g^{\wedge_{\zeta^{\prime}_{I}}}|_{L^{2}(d\eta^{\prime}_{I})}<C,\quad(1+{\eta^{\prime}_{I}}^{2})g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})\in L^{1}(d\eta^{\prime}_{I}).$
Besides, recall the Airy function
(3.45) $Ai(z)=\frac{1}{2\pi}\int_{\mathbb{R}}e^{i(\frac{s^{3}}{3}+zs)}ds$
which satisfies
(3.46) $|Ai(z)|\leq C(1+|z|)^{-\frac{1}{4}},\quad z\in\mathbb{R},$
(3.47) $Ai(-x)\sim\frac{1}{\sqrt{\pi}x^{\frac{1}{4}}}\cos\left(\frac{2}{3}x^{\frac{3}{2}}-\frac{\pi}{4}\right)+\mathcal{O}(x^{-\frac{7}{4}}),\quad x\to\infty,$
(3.48) $\left(e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}\right)^{\wedge_{\zeta^{\prime}_{I}}}(-\eta^{\prime}_{I})=\frac{2\pi}{(6t)^{\frac{1}{3}}}Ai\left(\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}(a-\frac{\pi\eta^{\prime}_{I}}{t})\right).$
Using (3.44), the Fourier multiplication formula, (3.46), and (3.48), (3.43) turns into
(3.49) $u_{1,1}(x)=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\eta^{\prime}_{I}\left(e^{-2it(a\zeta^{\prime}_{I}-{\zeta^{\prime}_{I}}^{3})}\right)^{\wedge_{\zeta^{\prime}_{I}}}(-\eta^{\prime}_{I})g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}})$
$=$ $-\frac{4}{(6t)^{\frac{1}{3}}}\frac{1}{\sqrt{t}}\int d\eta^{\prime}_{I}\ Ai\left(\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}(a-\frac{\pi\eta^{\prime}_{I}}{t})\right)g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}}).$
Moreover, let
(3.50) $z=\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}(a-\frac{\pi\eta^{\prime}_{I}}{t}),\quad\eta^{\prime}_{I}(t)=\frac{t}{\pi}(a+\frac{\sqrt[3]{3}}{(2t)^{\frac{2}{3}}}).$
Note that $\eta^{\prime}_{I}<-\frac{t}{C}r^{2}$ for $\eta^{\prime}_{I}<\eta^{\prime}_{I}(t)$ and $t\gg 1$. Hence from (3.44), as $t\to\infty$,
(3.51) $\begin{split}&|\theta(-\frac{t}{C}r^{2}-\eta^{\prime}_{I})g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})|_{L^{1}(d\eta^{\prime}_{I})}\leq Ct^{-1/2},\qquad|\frac{{\eta^{\prime}_{I}}^{2}}{t}\cdot g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})|_{L^{1}(d\eta^{\prime}_{I})}\leq Ct^{-1}.\end{split}$
Consequently, (3.49) implies
(3.52) $u_{1,1}(x)\leq$ $-\frac{4}{(6t)^{\frac{1}{3}}}\frac{1}{\sqrt{t}}\int_{\eta^{\prime}_{I}>\eta^{\prime}_{I}(t)}d\eta^{\prime}_{I}\ Ai\left(\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}(a-\frac{\pi\eta^{\prime}_{I}}{t})\right)g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(t^{-4/3}).$
Finally, for $\eta^{\prime}_{I}>\eta^{\prime}_{I}(t)$, we have z<1z<-1 and the Airy analysis (3.47) applies to (3.48). Along with the mean value theorem and (3.51), yields
$u_{1,1}(x)$
$=$ $\frac{-2}{(6t)^{\frac{1}{3}}\sqrt{t}}\int_{\eta^{\prime}_{I}>\eta^{\prime}_{I}(t)}d\eta^{\prime}_{I}\frac{e^{i(\frac{2}{3}\left|\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}|a-\frac{\pi\eta^{\prime}_{I}}{t}|\right|^{\frac{3}{2}}-\frac{\pi}{4})}+e^{-i(\frac{2}{3}\left|\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}|a-\frac{\pi\eta^{\prime}_{I}}{t}|\right|^{\frac{3}{2}}-\frac{\pi}{4})}}{\sqrt{\pi}\left[\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}|a-\frac{\pi\eta^{\prime}_{I}}{t}|\right]^{\frac{1}{4}}}g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{4}{3}}})$
$=$ $-\frac{2}{(6t)^{\frac{1}{2}}}\frac{1}{\sqrt{\pi rt}}\int_{\eta^{\prime}_{I}>\eta^{\prime}_{I}(t)}d\eta^{\prime}_{I}\left[e^{i(4tr^{3}(1-\frac{3}{2}\frac{\pi\eta^{\prime}_{I}}{t{\color[rgb]{0,0,0}|a|}})+\mathcal{O}(\frac{{\eta^{\prime}_{I}}^{2}}{t})-\frac{\pi}{4})}+c.c.\right]g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{4}{3}}})$
$=$ $\frac{2ie^{i4\pi tr^{3}}}{3t}\widetilde{s}_{c}(+ir)-\frac{2ie^{-i4\pi tr^{3}}}{3t}\widetilde{s}_{c}(-ir)+\mathcal{O}(\frac{1}{t^{\frac{4}{3}}})$
where c.c. denotes the complex conjugate of the preceding number. Therefore, we prove (3.39).
[…]
$u_{1,1}(x)=$ $-\frac{1}{3\pi t}\int d\zeta^{\prime}_{R}\ \int d\zeta^{\prime}_{I}\ e^{2\pi itS_{0}(a;\zeta^{\prime})}\psi_{r,0}(\zeta^{\prime}_{I})\partial_{\zeta^{\prime}_{R}}\left(\frac{1}{\zeta_{R}^{\prime}}\psi_{r,r}(\zeta^{\prime}_{R})\widetilde{s}_{c}(\zeta^{\prime})\right).$
Let $g_{+}=\psi_{r,0}(\zeta^{\prime}_{I})\partial_{\zeta^{\prime}_{R}}\left(\frac{1}{\zeta_{R}^{\prime}}\psi_{r,r}(\zeta^{\prime}_{R})\widetilde{s}_{c}(\zeta^{\prime})\right)$. Applying Lemma 3.3, and $u_{0}\in\mathfrak{M}^{1,q}$, we have
$|g_{+}|_{L^{2}(d\zeta^{\prime}_{I})},\ |\partial_{\zeta^{\prime}_{I}}g_{+}|_{L^{2}(d\zeta^{\prime}_{I})},\ |g_{+}^{\wedge_{\zeta^{\prime}_{I}}}|_{L^{1}(d\eta^{\prime}_{I})}<C.$
Therefore, taking the Fourier transform, and applying the Airy function analysis in the above proof, we obtain:
(3.53) $|u_{1,1}|\leq C|\frac{2\pi}{(6t)^{\frac{4}{3}}}\int d\zeta^{\prime}_{R}\int d\eta^{\prime}_{I}\,Ai\left(\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}(a-3{\zeta^{\prime}_{R}}^{2}-\frac{\pi\eta^{\prime}_{I}}{t})\right)g_{+}^{\wedge_{\zeta^{\prime}_{I}}}(\zeta_{R}^{\prime},\eta^{\prime}_{I})|\leq\frac{C.}{t^{\frac{4}{3}}}.$
We conclude this subsection by:
Theorem 3.
Suppose that (2.1) holds for $\mathfrak{M}^{2,q}$. Then, as $t\to+\infty$,
[…]
4. Long time asymptotics of $u_{2,0}(x)$
Throughout this subsection, $a$, $r$, $_{i}, tt$ are as defined in Definition 1 and the assumption of Theorem 1 holds. To study the asymptotics of $u_{2,0}$, it is necessary to analyze $(\widetilde{m}-1)$ and its derivatives. From
(4.1) $m=$ $1+\mathcal{C}T1+\cdots+\left(\mathcal{C}T\right)^{n}1+\cdots,$
we are led to study the Cauchy integrals $\widetilde{\left(\mathcal{C}T\right)^{n}1}$ and their derivatives.
4.1. The Cauchy integrals
4.1.1. Representation formulas of the Cauchy integrals
Lemma 4.1.
[8] Suppose the assumption of Theorem 1 holds.
$|\partial_{x_{1}}^{j}\widetilde{\mathcal{C}Tf}|_{L^{\infty}}\leq C|f|_{L^{\infty}},\quad j=0,\,1.$
Proof.
The proof follows from (2.12), (2.13), and
(4.2) $\partial_{x_{1}}^{j}\widetilde{\mathcal{C}Tf}$= $-\frac{1}{2\pi i}\iint\frac{e^{2\pi itS_{0}}(2\pi i\xi_{1}^{\prime})^{j}\widetilde{s}_{c}(\zeta^{\prime})}{\zeta^{\prime}-\lambda^{\prime}}\widetilde{f}(\zeta^{\prime})d\overline{\zeta}^{\prime}\wedge d\zeta^{\prime}$
${\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}\iint\frac{e^{2\pi itS_{0}(a;\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}(2\pi i\xi_{1}^{\prime})^{j}\widetilde{s}_{c}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi_{2}^{\prime})}\widetilde{f}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))d\xi^{\prime}_{1}d\xi^{\prime}_{2}},$
with
(4.3) $\begin{gathered}p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi^{\prime}_{2})=(2\pi i\xi^{\prime}_{1}+\lambda^{\prime})^{2}-(2\pi i\xi^{\prime}_{2}+{\lambda^{\prime}}^{2}),\\ \left|\frac{1}{p_{\lambda^{\prime}}}\right|_{L^{1}(\Omega_{\lambda^{\prime}},d\xi^{\prime}_{1}d\xi_{2}^{\prime})}\leq\frac{C}{(1+|\lambda^{\prime}_{I}|^{2})^{1/2}},\quad\left|\frac{1}{p_{\lambda^{\prime}}}\right|_{L^{2}(\Omega_{\lambda^{\prime}}^{c},d\xi^{\prime}_{1}d\xi^{\prime}_{2})}\leq\frac{C}{(1+|\lambda^{\prime}_{I}|^{2})^{1/4}},\\ |\xi_{1}^{j}s_{c}|_{L^{\infty}\cap L^{2}(d\xi_{1}d\xi_{2})}\sim|{\xi_{1}^{\prime}}^{j}\widetilde{s}_{c}|_{L^{\infty}\cap L^{2}(d\xi^{\prime}_{1}d\xi_{2}^{\prime})},\end{gathered}$
where $\Omega_{\lambda^{\prime}}=\{(\xi^{\prime}_{1},\xi^{\prime}_{2})\in{\mathbb{R}}^{2}:|p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi^{\prime}_{2})|<1\}$.
To study the long time asymptotics of the Cauchy integrals, inspired by [1] (cf. (3.9)), we present new representation formulas for $\widetilde{(\mathcal{C}T)^{n}1}$ in Lemma 4.2 and 4.4.
Lemma 4.2.
If the assumption of Theorem 1 holds then
(4.4) $\widetilde{\mathcal{C}T1}(t_{1},t_{2},t;\lambda^{\prime})=e^{i\pi tS_{0}(a;\lambda^{\prime})}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ [u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
$\quad\times\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{2}^{\prime};\xi_{1}^{\prime\prime})\equiv e^{i\pi tS_{0}(a;\lambda^{\prime})}[\mathfrak{C}\mathfrak{T}1]^{0,(1)}\equiv e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{0,(1)}1$
is holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when ${\lambda^{\prime}}_{I}\neq 0$. Here
(4.5) $\begin{gathered}\mathfrak{m}_{0}(x^{\prime}_{1},x^{\prime}_{2})-1=\iint(m_{0}(x_{1},x_{2};\overline{\zeta(\xi_{1},\xi_{2})})-1)^{\wedge_{x_{1},x_{2}}}e^{2\pi i(x^{\prime}_{1}\xi_{1}+x^{\prime}_{2}\xi_{2})}d\xi_{1}d\xi_{2},\\ |\partial_{x_{1}^{\prime}}^{j}(\mathfrak{m}_{0}-1)|_{L^{\infty}}\leq|\left(\partial^{j}_{x_{1}}(m_{0}(x_{1},x_{2};\overline{\zeta(\xi_{1},\xi_{2})})-1)\right)^{\wedge_{x_{1},x_{2}}}|_{L^{1}(d\xi_{1}d\xi_{2})}\leq C,\ \end{gathered}$
and $j=0,1$, with $\theta$ being the Heaviside function,
$e^{2\pi it\mathfrak{S}^{\sharp}(a,t;x_{1}^{\prime},x_{2}^{\prime};\lambda^{\prime}_{R};\xi_{1}^{\prime\prime})}=e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a-3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}+2{\lambda^{\prime}}_{R}x_{2}^{\prime}}{t})\xi_{1}^{\prime\prime}]}=e^{2\pi it\mathfrak{S}}e^{-2\pi i(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})\xi_{1}^{\prime\prime}},$
(4.6) $\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})=4\pi^{2}\xi_{1}^{\prime\prime 3}+(a-3{\lambda^{\prime}}_{R}^{2})\xi_{1}^{\prime\prime},$
$\mathcal{F}(t;\lambda^{\prime};x_{2}^{\prime};\xi_{1}^{\prime\prime})={\color[rgb]{0,0,0}\operatorname{sgn}(x_{2}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
$\hskip 108.405pt\times e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}.$
Proof.
Using (2.12), Lemma 4.1, the Fourier transform theory, $\exp{(+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}}))}$ is holomorphic in $\xi^{\prime}_{2}$ when $\xi_{1}^{\prime}\neq 0$ (i.e., holomorphic in $\zeta^{\prime}_{R}\zeta_{I}^{\prime}$ when $\zeta_{I}^{\prime}\neq 0$), and the residue theorem, we formally derive
(4.7) $\widetilde{\mathcal{C}T1}={\color[rgb]{0,0,0}\iint}[\,\frac{e^{+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})}}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi^{\prime}_{2})}\,]^{\vee_{\xi^{\prime}_{1},\xi^{\prime}_{2}}}(ta-x_{1}^{\prime},-x_{2}^{\prime})[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x_{2}^{\prime},x^{\prime}_{2})dx_{1}^{\prime}dx_{2}^{\prime},$
where $\mathfrak{m}_{0}$ satisfies (4.5) (see Lemma A.1 in the Appendix for the proof) and
(4.8) $[\ \frac{e^{+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})}}{p_{\lambda^{\prime}}({\xi^{\prime}_{1}},{\xi^{\prime}_{2}})}\ ]^{\vee_{{\xi^{\prime}_{1}},{\xi^{\prime}_{2}}}}(ta-x_{1}^{\prime},-x_{2}^{\prime})$
$=$ $-\frac{1}{2\pi i}\int d{\xi^{\prime}_{1}}\int d{\xi^{\prime}_{2}}\frac{e^{2\pi i[+t(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})+[(ta-x_{1}^{\prime}){\xi^{\prime}_{1}}-x_{2}^{\prime}{\xi^{\prime}_{2}}]]}}{{\xi^{\prime}_{2}}-(2\pi i{\xi^{\prime}_{1}}^{2}+2{\xi^{\prime}_{1}}\lambda^{\prime})}$
$\equiv$ $\frac{1}{2i}\int d{\xi^{\prime}_{1}}\ H_{2\pi i{\xi^{\prime}_{1}}^{2}+2{\xi^{\prime}_{1}}\lambda^{\prime}}(e^{2\pi i[+t(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})+[(ta-x_{1}^{\prime}){\xi^{\prime}_{1}}-x_{2}^{\prime}{\xi^{\prime}_{2}}]\ ]}).$
Here we have used (4.3), and
(4.9) $H_{s}(u)=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{u({\xi^{\prime}_{2}})}{s-{\xi^{\prime}_{2}}}d{\xi^{\prime}_{2}}$
which is holomorphic in $s\in\mathbf{C}^{\pm}$, and satisfies $H_{s^{+}}(u)-H_{s^{-}}(u)=-2iu(s)$ for $s\in{\mathbb{R}}$. Using the discontinuity is measure zero in $\xi^{\prime}_{1}$}, the residue theorem, $\xi_{1}^{\prime}=\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi}$,
$2\pi i\left[(ta-x_{1}^{\prime}){\xi^{\prime}_{1}}+t\pi^{2}{\xi^{\prime}_{1}}^{3}-x_{2}^{\prime}{\xi^{\prime}_{2}}-t\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}}\right]_{{\xi^{\prime}_{2}}=2\pi i{\xi^{\prime}_{1}}^{2}+2{\xi^{\prime}_{1}}\lambda^{\prime}}$
$=$ $2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a-3{\lambda^{\prime}}_{R}^{2})\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi}(a-3{\lambda^{\prime}}_{R}^{2}+{\lambda^{\prime}}_{I}^{2})]$
- $2\pi i(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})+4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}),$
and
$\operatorname{sgn}\left(\mathfrak{Im}(2\pi i{\xi^{\prime}_{1}}^{2}+2{\xi^{\prime}_{1}}\lambda^{\prime})\right)=\operatorname{sgn}((\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))=-\operatorname{sgn}(x_{2}^{\prime}+3t\lambda_{R}^{\prime})$
on the support of $\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$, we obtain
(4.10) $\left[\frac{e^{-2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})}}{p_{\lambda^{\prime}}({\xi^{\prime}_{1}},{\xi^{\prime}_{2}})}\right]^{\vee_{{\xi^{\prime}_{1}},{\xi^{\prime}_{2}}}}(ta-x_{1}^{\prime},-x_{2}^{\prime})=\operatorname{sgn}(x_{2}^{\prime}+3t\lambda_{R}^{\prime})$
$\times$ $e^{-it(a\lambda^{\prime}_{I}+{\lambda^{\prime}}_{I}^{3}-3\lambda^{\prime}_{I}{\lambda^{\prime}}_{R}^{2})}e^{-2\pi i(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})(-\frac{\lambda^{\prime}_{I}}{2\pi})}\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a-3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime}}{t})\xi_{1}^{\prime\prime}]}$
$\times$ $\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}.$
Plugging (4.10) into (4.7), we justify (4.4), (4.6), and holomorphicy in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when $\lambda^{\prime}_{I}\neq 0$ formally.
For the rigorous analysis, we first show the uniform boundedness when $\mathcal{F}$ fails to decay :
(4.11) $C\geq$ $\lim_{x_{2}^{\prime}+3t\lambda^{\prime}_{R}\to 0^{\pm}}\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a-3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime}}{t})\xi_{1}^{\prime\prime}]}$
$\times$ $\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}$
$=$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})\xi_{1}^{\prime\prime}]}\theta((a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})-1)$
$+$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})\xi_{1}^{\prime\prime}]}\theta(1-|a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t}|)$
$+$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})\xi_{1}^{\prime\prime}]}\theta(-1-(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t}))\equiv I+II+III.$
Integration by parts, using $(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})>1$, we obtain $|I|\leq C$. Similarly,
(4.12) $|II|\leq$ $|\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})\xi_{1}^{\prime\prime}]}\theta(1-|a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t}|)\theta(1-|\xi_{1}^{\prime\prime}|)$
$+$ $|\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})\xi_{1}^{\prime\prime}]}\theta(1-|a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t}|)\theta(|\xi_{1}^{\prime\prime}|-1)\leq C,$
(4.13) $|III|\leq$ $|\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})\xi_{1}^{\prime\prime}]}\theta(-1-(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t}))\psi_{1,\rho}(\xi_{1}^{\prime\prime})$
$+$ $|\int d\xi_{1}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{1}^{\prime\prime 3}+(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})\xi_{1}^{\prime\prime}]}\theta(-1-(a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t}))(1-\psi_{1,\rho}(\xi_{1}^{\prime\prime}))\leq C$
by letting $\pm\rho=\pm\left[|a+3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1}^{\prime}}{t})|\right]^{1/2}$, using integration by parts and $|\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}^{\sharp}|>1/C$ for the second terms. Combining II-IIIIII, the uniform boundedness of (4.11) is proved.
Therefore, assuming $u_{0}\in{\mathfrak{M}^{0,q}}$, and using the estimate (4.5) (see Lemma A.1 in the Appendix), the representation formula (4.4) holds rigorously and is holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when $\lambda^{\prime}_{I}\neq 0$.
To apply an inductive argument to derive the representation formulas for $\widetilde{(\mathcal{C}T)^{n}1}$, particularly in generalizing the reasoning used in (4.11), we require:
Lemma 4.3.
If the assumption of Theorem 1 holds for $u_{0}\in{\mathfrak{M}^{1,q}}$, then we have:
(4.14) $|\partial_{\lambda^{\prime}_{I}}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(1)}|\leq C(1+|\lambda_{R}^{\prime}|).$
Proof.
From (4.4),
(4.15) $|\partial_{\lambda^{\prime}_{I}}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(1)}|_{L^{\infty}}$
$\leq$ $C|\iint dx_{1}^{\prime}dx_{2}^{\prime}\ (x_{1}^{\prime}+2x_{2}^{\prime}\lambda^{\prime}_{R})[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}|$
$+$ $C\iint dx_{1}^{\prime}dx_{2}^{\prime}\ |[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})|$
$+$ $C\iint dx_{1}^{\prime}dx_{2}^{\prime}|[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})||\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\theta(-(3t\lambda^{\prime}_{R}+x_{2}^{\prime})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
$\times$ $\lambda^{\prime}_{I}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})e^{\ 4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}|\equiv I_{1}+I_{2}+I_{3}.$
From $\color[rgb]{0,0,0}u_{0}\in\mathfrak{M}^{0,q}$ and Lemma 4.1,
(4.16) $I_{2}\leq C.$
Theorem 2 and $u_{0}\in\mathfrak{M}^{1,q}$ imply that there exist $s_{\sharp,k}$, $u_{\sharp,k}$, $m_{\sharp,k}$, $k=1,2$, such that
(4.17) $\begin{gathered}x_{k}^{\prime}[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})=[u_{\sharp,k}\mathfrak{m}_{\sharp,k}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2}),\end{gathered}$
and
(4.18) $u_{\sharp,k}(x_{1},x_{2})=\overline{u_{\sharp,k}(x_{1},x_{2})},\quad u_{\sharp,k}\in{\color[rgb]{0,0,0}\mathfrak{M}^{0,q-4}},\quad|u_{\sharp,k}|_{\mathfrak{M}^{0,0}}\ll 1,$
$m_{\sharp,k}$ satisfy (2.7) with $T$ replaced by $T_{\sharp,k}$ and $x_{3}=0$;
$T_{\sharp,k}$ defined by (2.9) with $s_{c}$ replaced by $s_{\sharp,k}=\mathcal{S}(u_{\sharp,k})$;
[…]
Along with Lemma 4.1, Proposition 4.1, yield:
(4.19) $I_{1}\leq C|\mathcal{C}T_{\sharp,1}1|+C|\lambda^{\prime}_{R}||\mathcal{C}T_{\sharp,2}1|\leq C(1+|\lambda_{R}^{\prime}|).$
Finally, for $I_{3}$, notice
(4.20) ${\color[rgb]{0,0,0}\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))\lambda^{\prime}_{I}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}}$
(4.23) $\sim$ $\left\{{\begin{array}[]{ll}\theta(-\lambda^{\prime}_{I}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))\partial_{\xi_{1}^{\prime\prime}}e^{4\pi^{2}\lambda^{\prime}_{I}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})},&\textit{if $\xi_{1}^{\prime\prime}\lambda^{\prime}_{I}<0$;}\\ \theta(-\lambda^{\prime}_{I}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi}))\partial_{\xi_{1}^{\prime\prime}}e^{4\pi^{2}\lambda^{\prime}_{I}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})},&\textit{if $\xi_{1}^{\prime\prime}\lambda^{\prime}_{I}>0$. }\end{array}}\right$.
Hence
(4.24) $I_{3}\leq C.$
Lemma 4.4.
If the assumption of Theorem 1 holds for $u_{0}\in{\mathfrak{M}^{1,q}}$ and $\color[rgb]{0,0,0}n\geq 2$, then
(4.25) $\widetilde{(\mathcal{C}T)^{n}1}(t_{1},t_{2},t;\lambda^{\prime})=e^{\beta_{n}i\pi tS_{0}(a;\lambda^{\prime})}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
where
$\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n)}$ $(t_{1},t_{2},t;\lambda^{\prime})=\iint dx_{1,n}^{\prime}dx_{2,n}^{\prime}[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})e^{\beta_{n}i\lambda^{\prime}_{I}(x_{1,n}^{\prime}+2\lambda^{\prime}_{R}x_{2,n}^{\prime})}$
$\times$ $\int d\xi_{n}^{\prime\prime}e^{\beta_{n}2\pi it\mathfrak{G}^{\sharp}(a,t;x^{\prime}_{1,n},x^{\prime}_{2,n};\lambda_{R}^{\prime};\xi_{n}^{\prime\prime})}\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{n}^{\prime\prime})$
$\equiv$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{n}^{\prime\prime}),$
and
$\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(0)}=1,\ x_{1,1}^{\prime}=x_{1}^{\prime},\ x_{2,1}^{\prime}=x_{2}^{\prime},$
$\frac{1}{2}\leq\beta_{n}=\frac{1}{2}(2-\beta_{n-1})\leq 1,\ \beta_{1}=1,$
$\mathcal{F}^{(n)}(t;\lambda^{\prime};x_{2,n}^{\prime};\xi_{n}^{\prime\prime})={\color[rgb]{0,0,0}\operatorname{sgn}(x_{2,n}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
$\times e^{\beta_{n}4\pi^{2}(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}.$
Moreover, $\widetilde{(\mathcal{C}T)^{n}1}$ is holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when ${\lambda^{\prime}}_{I}\neq 0$, and
(4.26) $|\partial_{\lambda^{\prime}_{I}}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n)}|\leq C(1+|\lambda_{R}^{\prime}|).$
Proof.
Once (4.25) is established, the proof of (4.26) can be established using the same argument as that for Lemma 4.3. Hence it is sufficient to justify (4.25).
Using Lemma 4.1, 4.2,
$\overline{\zeta^{\prime}}|_{\xi_{2}^{\prime}=2\pi i{\xi_{1}^{\prime}}^{2}+2\xi_{1}^{\prime}\lambda^{\prime},\ \xi_{1}^{\prime}=\xi_{n}^{\prime\prime}-\frac{\lambda_{I}^{\prime}}{2\pi}}=\lambda_{R}^{\prime}+2\pi i\xi_{n}^{\prime\prime},$
and an induction, formally we obtain:
$\widetilde{(\mathcal{C}T)^{n}1}(t_{1},t_{2},t;\lambda^{\prime})$
$=$ ${\color[rgb]{0,0,0}-\frac{1}{2\pi i}}\iint\left[\frac{e^{\frac{2-\beta_{n-1}}{2}2\pi iS_{0}(\zeta^{\prime})}\left[\mathfrak{C}\mathfrak{T}\right]^{(n-1)}(t_{1},t_{2},t;\overline{\zeta}^{\prime})}{p_{\lambda^{\prime}}({\xi^{\prime}_{1}},{\xi^{\prime}_{2}})}\right]^{\vee_{{\xi^{\prime}_{1}},{\xi^{\prime}_{2}}}}(ta-x_{1,n}^{\prime},-x_{2,n}^{\prime})$
$\times$ $[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})dx_{1,n}^{\prime}dx_{2,n}^{\prime}$
$=$ $e^{\beta_{n}i\pi tS_{0}(\lambda^{\prime};a)}\iint dx_{1,n}^{\prime}dx_{2,n}^{\prime}[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})e^{\beta_{n}i\lambda^{\prime}_{I}(x_{1,n}^{\prime}+2\lambda^{\prime}_{R}x_{2,n}^{\prime})}$
$\times$ $\int d\xi_{n}^{\prime\prime}\ e^{\beta_{n}2\pi it\mathfrak{G}^{\sharp}(a,t;x^{\prime}_{1,n},x^{\prime}_{2,n};\lambda_{R}^{\prime};\xi_{n}^{\prime\prime})}\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{n}^{\prime\prime})$
$=$ $e^{\beta_{n}i\pi tS_{0}(\lambda^{\prime};a)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n)}(t_{1},t_{2},t;\lambda^{\prime}).$
To make the above formula hold rigorously, be holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when ${\lambda^{\prime}}_{I}\neq 0$, beyond the argument in Lemma 4.2, the key step here is to justify the uniformly boundedness of corresponding (4.11) using integration by parts. Precisely,
(4.27) $\lim_{x_{2,n}^{\prime}+3t\lambda^{\prime}_{R}\to 0^{\pm}}\int d\xi_{n}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{n}^{\prime\prime 3}+(a-3{\lambda^{\prime}}_{R}^{2}-\frac{x_{1,n}^{\prime}+2\lambda^{\prime}_{R}x_{2,n}^{\prime}}{t})\xi_{n}^{\prime\prime}]}\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(\lambda^{\prime}_{R}+2\pi i\xi_{n}^{\prime\prime})$
$=$ $\int d\xi_{n}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{n}^{\prime\prime 3}+(a+3{\lambda^{\prime}_{R}}^{2}-\frac{x_{1,n}^{\prime}}{t})\xi_{n}^{\prime\prime}]}\theta((a+3{\lambda^{\prime}_{R}}^{2}-\frac{x_{1,n}^{\prime}}{t})-1)$
$\times$ $\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(t_{1},t_{2},t;\lambda^{\prime}_{R}+2\pi i\xi_{n}^{\prime\prime})$
$+$ $\int d\xi_{n}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{h}^{\prime\prime 3}+(a+3{\lambda^{\prime}_{R}}^{2}-\frac{x_{1,n}^{\prime}}{t})\xi_{n}^{\prime\prime}]}\theta(1-|a+3{\lambda^{\prime}_{R}}^{2}-\frac{x_{1,n}^{\prime}}{t}|)$
$\times$ $\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(t_{1},t_{2},t;\lambda^{\prime}_{R}+2\pi i\xi_{n}^{\prime\prime})$
$+$ $\int d\xi_{n}^{\prime\prime}e^{2\pi it[4\pi^{2}\xi_{n}^{\prime\prime 3}+(a+3{\lambda^{\prime}_{R}}^{2}-\frac{x_{1,n}^{\prime}}{t})\xi_{n}^{\prime\prime}]}\theta(-1-(a+3{\lambda^{\prime}_{R}}^{2}-\frac{x_{1,n}^{\prime}}{t}))$
$\times$ $\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(t_{1},t_{2},t;\lambda^{\prime}_{R}+2\pi i\xi_{n}^{\prime\prime})\equiv I^{(n)}+II^{(n)}+III^{(n)}.$
Integration by parts, using Lemma 4.3, and (4.26) inductively, analogous to Lemma 4.2,
(4.28) $|I^{(n)}|,\,|II^{(n)}|,\,|III^{(n)}|\leq C(1+|\lambda_{R}^{\prime}|).$
Thanks to $u_{0}\in{\mathfrak{M}^{1,q}}$, we have
$\lim_{x_{2,n}^{\prime}+3t\lambda^{\prime}_{R}\to 0^{\pm}}|(1+|\lambda^{\prime}_{R}|)[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})|$
$\leq|(1+|x_{1,n}^{\prime}|+|x_{2,n}^{\prime}|)[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})|.$
Hence, for $u_{0}\in\mathfrak{M}^{1,q}$, the representation formula (4.25) holds rigorously and is holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when ${\lambda^{\prime}}_{I}\neq 0$.
Definition 2.
Let the phase function $\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})$ be defined by (4.6). In view of
(4.29) $\begin{split}\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})=&+12\pi^{2}{\xi_{1}^{\prime\prime}}^{2}+(a-3{\lambda^{\prime}}_{R}^{2}),\\ \partial_{\xi_{1}^{\prime\prime}}^{2}\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})=&+24\pi^{2}{\xi_{1}^{\prime\prime}},\end{split}$
we have the definition for stationary points :
If $a-3{\lambda^{\prime}}_{R}^{2}>0$, there are no stationary points of $\mathfrak{S}$.
If $a-3{\lambda^{\prime}}_{R}^{2}<0$, there are two stationary points ${\xi_{1}^{\prime\prime}}=\pm b\gtrless 0$, $b^{2}=\frac{3{\lambda^{\prime}}_{R}^{2}-a}{12\pi^{2}}$, of $\mathfrak{S}$.
4.1.2. Asymptotics of the Cauchy integrals
Proposition 4.1.
If the assumption of Theorem 1 holds for $u_{0}\in{\mathfrak{M}^{1,q}}$ then
(4.30) $|\widetilde{\left(\mathcal{C}T\right)^{n}1}|\leq\mathcal{O}(t^{-1/2})\quad\textit{as $t\to\infty$}.$
Proof.
Applying Lemmas 4.1, 4.2, and 4.4, it reduces to studying the asymptotics of $\mathfrak{C}\mathfrak{T}1$.
Denote
(4.31) $\Sigma_{s}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})=\{\xi_{1}^{\prime\prime}:|\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})|\leq s\}.$
From Definition 2, we obtain estimates for the measure $|\Sigma_{s}|$,
(4.32) $|\Sigma_{t^{-1/2}}|\leq Ct^{-1/2}.$
Hence, if $|\lambda^{\prime}_{R}|<r$, then using integration by parts, (4.32), and $u_{0}\in\mathfrak{M}^{1,q}$, we get
(4.33) $|\mathfrak{C}\mathfrak{T}1|\leq$ $|\mathfrak{C}\mathfrak{T}\theta(t^{-1/2}-|\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})|)|+|\mathfrak{C}\mathfrak{T}\theta(|\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})|-t^{-1/2}))|$
$\leq$ $Ct^{-1/2}+\frac{C}{t}|\iint dx_{1}^{\prime}dx_{2}^{\prime}\ [u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
$\times$ $\int_{\xi_{1}^{\prime\prime}\in\Sigma_{t^{-1/2}}^{c}}d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}}\partial_{\xi_{1}^{\prime\prime}}\frac{1}{\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}}\{e^{-2\pi i(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})\xi_{1}^{\prime\prime}}\operatorname{sgn}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})$
$\times$ $\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}\}|\leq C{t^{-1/2}}.$
If $|\lambda^{\prime}_{R}|>r$, using $u_{0}\in\mathfrak{M}^{1,q}$, we obtain
(4.34) $|\mathfrak{C}\mathfrak{T}1|\leq|\mathfrak{C}\mathfrak{T}\theta(t|\lambda^{\prime}_{R}|-|x_{2}^{\prime}|)|+\epsilon_{0}\mathcal{O}(t^{-1})$
$\ \leq Ct^{-1/2}+\frac{C}{t}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ |[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})\theta(t|\lambda^{\prime}_{R}|-|x_{2}^{\prime}|)x_{2}^{\prime}|$
$\ \times|\int_{\xi_{1}^{\prime\prime}\in\Sigma_{t^{-1/2}}^{c}}d{\xi_{1}^{\prime\prime}}\frac{1}{|\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}|}\psi_{r,0}(\xi_{1}^{\prime\prime})\cdot\lambda_{R}^{\prime}\cdot\mathcal{F}|$
$\ +\frac{C}{t^{2}}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ |[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})\theta(t|\lambda^{\prime}_{R}|-|x_{2}^{\prime}|)x_{2}^{\prime}|$
$\ \times|\int_{\xi_{1}^{\prime\prime}\in\Sigma_{t^{-1/2}}^{c}}d{\xi_{1}^{\prime\prime}}\frac{1}{\partial_{\xi_{1}^{\prime\prime}}\mathfrak{S}}[1-\psi_{r,0}(\xi_{1}^{\prime\prime})]\cdot t\lambda_{R}^{\prime}\xi_{1}^{\prime\prime}\cdot\mathcal{F}|\equiv Ct^{-1/2}+I_{1}+I_{2}.$
Notice that if $\xi_{1}^{\prime\prime}\in\{\xi_{1}^{\prime\prime}:\xi_{1}^{\prime\prime}\in\Sigma_{t^{-1/2}}^{c}\}\cap\{\xi_{1}^{\prime\prime}:\psi_{r,0}(\xi_{1}^{\prime\prime})\neq 0\}$, then
${\lambda_{R}^{\prime}}^{2}\leq 4{\color[rgb]{0,0,0}\pi r}^{2}+\frac{a}{3}{\color[rgb]{0,0,0}+}\frac{1}{t^{1/2}}.$
Consequently,
(4.35) $|I_{1}|\leq C{t^{-1/2}}.$
On the other hand,
(4.36) $|I_{2}|\leq C{t^{-1/2}}$
follows from $u_{0}\in\mathfrak{M}^{1,q}$ and
$\theta(t|\lambda^{\prime}_{R}|-|x_{2}^{\prime}|)t\lambda_{R}^{\prime}\xi_{1}^{\prime\prime}\cdot\mathcal{F}\sim\theta(t|\lambda^{\prime}_{R}|-|x_{2}^{\prime}|)\\ \times\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))\partial_{\xi_{1}^{\prime\prime}}e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}.$
Hence
(4.37) $|\mathfrak{C}\mathfrak{T}1|\leq C{t^{-1/2}},\quad\textit{ for $|\lambda^{\prime}_{R}|>r$}.$
Applying Lemma 4.1-4.4, and Proposition 4.1, we obtain the first reduction:
Proposition 4.2.
If the assumption of Theorem 1 holds for $u_{0}\in{\mathfrak{M}^{3,q}}$. For $\color[rgb]{0,0,0}|a|>\delta>0$, as $t\to\infty$,
(4.38) $|u_{2,0}(x)|\leq$ $C\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\widetilde{s}_{c}(\lambda^{\prime})e^{\beta_{n+1}2\pi itS_{0}}(\overline{\lambda}^{\prime}-\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|+o(t^{-1}).$
Proof.
If $\color[rgb]{0,0,0}|a|>\delta>0$, from (2.3), Lemma 4.2, 4.4, and Proposition 4.1, for $u_{0}\in\mathfrak{M}^{1,q}$, we obtain
(4.39) $|u_{2,0}(x)|\leq$ $C\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\widetilde{s}_{c}(\lambda^{\prime})e^{\beta_{n+1}2\pi itS_{0}}(\overline{\lambda}^{\prime}-\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|+o(t^{-1}).$
Besides, $u_{0}\in{\mathfrak{M}^{3,q}}$ implies
(4.40) $|\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(|x_{2,n}^{\prime}|-t|\lambda^{\prime}_{R}|)[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})|_{L^{1}(dx^{\prime}_{1,n}dx^{\prime}_{2,n})}$
$\leq$ $C\frac{|\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(|x_{2,n}^{\prime}|-t|\lambda^{\prime}_{R}|)(1+|x^{\prime}_{2,n}|)^{3}[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})|_{L^{1}(dx^{\prime}_{1,n}dx^{\prime}_{2,n})}}{(1+|t\lambda^{\prime}_{R}|)^{3}}$
$\leq$ $Ct^{-4/9\times 3}|u_{0}|_{{\mathfrak{M}^{3,q}}}.$
Along with (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, (4.39), and Lemma 4.2, 4.4, yields (4.38). ∎
Lemma 4.5.
If the assumption of Theorem 1 holds for $u_{0}\in{\mathfrak{M}^{1,q}}$ and $C_{0}>0$. As $t\to\infty$,
(4.41) $|\partial_{\lambda^{\prime}_{I}}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n)}|\leq$ $C(1+|\lambda_{R}^{\prime}|),$
(4.42) $\theta(|\lambda_{R}^{\prime}|-1/C_{0})|\partial_{\lambda^{\prime}_{R}}[\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)[\mathfrak{C}\mathfrak{T}1]^{0,(n-1)}]\leq$ $C(1+|\lambda_{I}^{\prime}|).$
Proof.
Proof of (4.41) follows from the same argument used in the proof of Lemma 4.3. To prove (4.42), from (4.25),
$\partial_{\lambda^{\prime}_{R}}[\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)[\mathfrak{C}\mathfrak{T}1]^{0,(n-1)}]$
$\leq$ $C|\iint dx_{1,n}^{\prime}dx_{2,n}^{\prime}\ x_{2,n}^{\prime}\lambda_{I}^{\prime}[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})e^{\beta_{n}i\lambda_{I}^{\prime}(x_{1,n}^{\prime}+2\lambda^{\prime}_{R}x^{\prime}_{2,n})}$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\int d\xi_{n}^{\prime\prime}e^{\beta_{n}2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{n}^{\prime\prime})|$
$+$ ${\color[rgb]{0,0,0}C|\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|_{L^{\infty}}\int dx_{1,n}^{\prime}\ |t[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}\pm 2t_{2}t\lambda^{\prime}_{R},\pm t\lambda_{R}^{\prime})|}$
$+$ $C|\iint dx_{1,n}^{\prime}dx_{2,n}^{\prime}\ [u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})e^{\beta_{n}i\lambda_{I}^{\prime}(x_{1,n}^{\prime}+2\lambda^{\prime}_{R}x^{\prime}_{2,n})}$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\int d\xi_{n}^{\prime\prime}e^{\beta_{n}2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\cdot 4\pi i(x_{2,h}^{\prime}+3t\lambda^{\prime}_{R})\xi_{h}^{\prime\prime}|$
$+$ $C|\iint dx_{1,n}^{\prime}dx_{2,n}^{\prime}\ [u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})e^{\beta_{n}i\lambda_{I}^{\prime}(x_{1,n}^{\prime}+2\lambda^{\prime}_{R}x^{\prime}_{2,n})}$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\int d\xi_{n}^{\prime\prime}e^{\beta_{n}2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\cdot t(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})(\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime})|$
$+$ $C|\iint dx_{1,n}^{\prime}dx_{2,n}^{\prime}\ [u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})e^{\beta_{n}i\lambda_{I}^{\prime}(x_{1,n}^{\prime}+2\lambda^{\prime}_{R}x^{\prime}_{2,n})}$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\int d\xi_{n}^{\prime\prime}e^{\beta_{n}2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}^{(n)}\partial_{\lambda^{\prime}_{R}}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
$\equiv$ ${\color[rgb]{0,0,0}I^{(n)}_{1}+I^{(n)}_{2}+I^{(n)}_{3}+I^{(n)}_{4}+I^{(n)}_{5}}$.
Applying $u_{0}\in{\mathfrak{M}^{1,q}}$, (4.18), and Lemma 4.1, we obtain
(4.43) $I^{(n)}_{1}\leq C|\lambda_{I}^{\prime}||\mathcal{C}T_{\sharp,2}(\mathcal{C}T)^{n-1}1|\leq C|\lambda_{I}^{\prime}|.$
Besides, for $\color[rgb]{0,0,0}u_{0}\in\mathfrak{M}^{1,q}$,
(4.44) $I^{(n)}_{3}\leq$ $C\iint dx_{1}^{\prime}dx_{2}^{\prime}|[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})|\,|\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}$
$\times$ $\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))\partial_{\xi_{1}^{\prime\prime}}e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}|\leq C,$
and
(4.45) $\theta(|\lambda_{R}^{\prime}|-1/C_{0})I^{(n)}_{2}\leq C,$
(4.46) $\theta(|\lambda_{R}^{\prime}|-1/C_{0})I^{(n)}_{4}\leq C\frac{\theta(|\lambda_{R}^{\prime}|-1/C_{0})}{|\lambda^{\prime}_{R}|}\iint dx_{1,n}^{\prime}dx_{2,n}^{\prime}\ |[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})|$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)|\int d\xi_{n}^{\prime\prime}e^{\beta_{n}2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}^{(n)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\xi_{n+1}^{\prime\prime})(\xi_{n}^{\prime\prime}+\xi_{n+1}^{\prime\prime})|$
$\leq$ $C.$
Applying Lemma 4.1, $u_{0}\in\mathfrak{M}^{1,q}$, and an induction, we obtain
(4.47) $I^{(n)}_{5}\leq C(1+|\lambda_{I}^{\prime}|).$
The above lemma shows that taking the derivatives of the Cauchy integrals, no matter how small a neighborhood is chosen around these points, the $(1+|\lambda^{\prime}|)\mathcal{O}(1)$ bounds on the $\lambda^{\prime}$-derivatives of the Cauchy integrals cannot be improved. Moreover, higher derivatives of the Cauchy integrals correspond to higher orders in $t$. This presents a fundamental obstruction to obtaining $o(t^{-1})$ estimates for $u_{2,0}$ and $u_{2,1}$ through our approach.
4.2. Long time asymptotics of $u_{2,0}(x)$ when $\color[rgb]{0,0,0}a>+\delta>0$
Throughout this subsection, the assumption of Theorem 1 holds for $\color[rgb]{0,0,0}a>+\delta>0$, define $\psi_{r,w_{0}}$, $\color[rgb]{0,0,0}\chi$ as in (3.20), (3.21), and set $b=(-r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$.
Building on Proposition 4.2 and Lemma 4.5, we can further focus our attention on the neighborhoods of the stationary points, as stated in the following lemma:
Lemma 4.6.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(4.48) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}[1-\chi(\lambda^{\prime})]}$
$\times\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|\leq o(t^{-1}).$
Proof.
The proof of the lemma demonstrates that the term $(\overline{\lambda^{\prime}}-\lambda^{\prime})$ is essential for eliminating the Kiselev conditions, such as the integrability of $(1+|\lambda^{\prime}|)s_{c}$ or boundedness of $\partial^{k}_{\lambda^{\prime}_{I}}s_{c}$ for $k\leq 2$.
[…]
    Case of $a-3{\lambda^{\prime}_{R}}^{2}>0$: In this case, $|{\lambda_{R}^{\prime}}|\leq r$, and the analysis can be reduced to cases:
    • (1)
      $\psi_{r,r}(\lambda^{\prime}_{R})\neq 0$ and $\psi_{r,0}(\lambda^{\prime}_{I})=0$;
    • (2)
      $\psi_{r,r}(\lambda^{\prime}_{R})=0$.
    Notice that $\partial_{\lambda^{\prime}_{I}}\mathfrak{S}(a;\lambda_{R}^{\prime};\lambda^{\prime}_{I})=+12\pi^{2}{\lambda^{\prime}_{I}}^{2}+(a-3{\lambda^{\prime}}_{R}^{2})\geq r/C$ for both cases. Therefore, we obtain (4.48) by applying integration by parts with respect to $\lambda^{\prime}_{I}$ to
    $\hskip 36.135pt{\color[rgb]{0,0,0}\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})[1-\chi(\lambda^{\prime})]\theta(a-3{\lambda^{\prime}_{R}}^{2})}\\ {\color[rgb]{0,0,0}\times\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)},}\hskip 36.135pt$
    • using (2.3), and $|{\lambda_{R}^{\prime}}|\leq r$ to obtain
      (4.49) $|\partial_{\lambda^{\prime}_{I}}\left(\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\frac{1}{|\partial_{\lambda^{\prime}_{I}}\mathfrak{S}|}\right)|_{L^{1}(d\lambda^{\prime}_{R}d\lambda^{\prime}_{I})}<C;$
    • adopting (2.3), (4.26), Lemma 3.3, and $|{\lambda_{R}^{\prime}}|\leq r$ to get
      (4.50) $|\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\frac{1}{|\partial_{\lambda^{\prime}_{I}}\mathfrak{S}|}\partial_{\lambda^{\prime}_{I}}\left[\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\right]|_{L^{1}(d\lambda^{\prime}_{R}d\lambda^{\prime}_{I})}<C,$
    and adapting the argument from the proof of Proposition 3.1.
[…]
    Proof of $a-3{\lambda^{\prime}_{R}}^{2}<0$: In this case, $|{\lambda_{R}^{\prime}}|\geq r$. The analysis can be reduced to cases:
    • (1’)
      $\psi_{r,r}(\lambda^{\prime}_{R})\neq 0$ and $\psi_{r,0}(\lambda^{\prime}_{I})=0$;
    • (2’)
      $\psi_{r,r}(\lambda^{\prime}_{R})=0$.
    Proof of Case of (1’) can be proceeded as that of Case (1). For Case of (2’), (4.48) is justified by adapting argument of Proposition 3.1, that is, integration by parts with respect to $\lambda^{\prime}$ to
    $\hskip 36.135pt{\color[rgb]{0,0,0}\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})[1-\chi(\lambda^{\prime})]\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}\\ {\color[rgb]{0,0,0}\times\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}},\hskip 36.135pt$
    • Using $\psi_{r,r}(\lambda^{\prime}_{R})=0$ and $|{\lambda_{R}^{\prime}}|\geq r$, we derive
      (4.51) $\nabla\cdot{\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}=0,\quad\nabla\cdot{\color[rgb]{0,0,0}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})}=0.$
    • Applying (2.3), Lemma 3.3, and taking advantage of the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, we have
      (4.52) $|\nabla\cdot\left(\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)|_{L^{1}(d\lambda^{\prime}_{R}d\lambda^{\prime}_{I})}<C.$
    • Applying $u_{0}\in\mathfrak{M}^{3,q}$, (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, and Lemma 4.5 for $|{\lambda_{R}^{\prime}}|\geq r$,
      (4.53) $|\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\nabla\cdot{\color[rgb]{0,0,0}\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}}|_{L^{1}(d\lambda^{\prime}_{R}d\lambda^{\prime}_{I})}<C.$
Lemma 4.7.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(4.54) $\hskip 36.135pt\sum_{n=1}^{\infty}\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\chi(\lambda^{\prime})}\\ \times\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\sim{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.\hskip 36.135pt$
Proof.
Notice
LHS of (4.54)
$=$ ${\color[rgb]{0,0,0}\frac{1}{3\pi t}\sum_{n=1}^{\infty}\int d\lambda^{\prime}_{I}\int d\lambda^{\prime}_{R}\ e^{\beta_{n+1}2\pi itS_{0}}\partial_{\lambda_{R}^{\prime}}\left(\frac{1}{\lambda^{\prime}_{R}}\widetilde{s}_{c}(\lambda^{\prime})\chi(\lambda^{\prime})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\right)}.$
Along with Lemmas 3.3, 4.4, and the cutoff function $\chi(\lambda^{\prime})$, yields (4.54).
Theorem 4.
Assume the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,
(4.55) $u_{2,0}\sim{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.$
Proof.
Follows from Proposition 4.2, Lemmas 4.6, and 4.7. ∎
4.3. Long time asymptotics of $u_{2,0}(x)$ when $\color[rgb]{0,0,0}a<-\delta<0$
Throughout this subsection, we assume the hypotheses of Theorem 1 and define $\psi_{r,w_{0}}$ as in (3.20). We also set $b=(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$ and adopt the terminology introduced in Lemmas 4.2 and 4.4.
An analogue of Lemma 4.6 is stated as follows:
Lemma 4.8.
Suppose the assumption of Theorem 1 is valid for $u_{0}\in\mathfrak{M}^{3,q}$. Then there exists a constant $C_{0}\gg 1$ such that, defining
(4.56) $\color[rgb]{0,0,0}\Xi(\lambda^{\prime})=\theta(C_{0}-|\lambda^{\prime}_{R}|)+\theta(|\lambda^{\prime}_{R}|-C_{0})\theta(|\lambda^{\prime}_{I}|-1),$}
we have
(4.57) $\color[rgb]{0,0,0}|\Xi(\lambda^{\prime})\widetilde{s}_{c}(\lambda^{\prime})|_{L^{1}(d\lambda^{\prime}_{R}d\lambda^{\prime}_{I})}\leq C,$}
and, as $t\to\infty$,
(4.58) ${\color[rgb]{0,0,0}|u_{2,0}(x)|\leq}$ ${\color[rgb]{0,0,0}\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\Xi(\lambda^{\prime})\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}$
$\times$ ${\color[rgb]{0,0,0}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|+\mathcal{O}(t^{-1}).}$
Proof.
Estimate (4.57) follows from (2.2) and (2.12) directly.
[…]
(4.59) $|u_{2,0}(x)|\leq$ $C\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\Xi(\lambda^{\prime})$
$\times$ $\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|+\mathcal{O}(t^{-1}),$
which reduces to showing
(4.60) $\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\theta(|\lambda^{\prime}_{R}|-C_{0})\theta(1-|\lambda^{\prime}_{I}|)$
$\times$ $\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\leq O(t^{-1}),$
by applying Proposition 4.2.
To this aim, we decompose the LHS of (4.60) into
(4.61) $\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\theta(|\lambda^{\prime}_{R}|-C_{0})\theta(1-|\lambda^{\prime}_{I}|)$
$\times$ $\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda_{I}^{\prime}|}{2\pi}|-1)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}$
$+$ $\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\theta(|\lambda^{\prime}_{R}|-C_{0})\theta(1-|\lambda^{\prime}_{I}|)$
$\times$ $\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(1-||\xi_{n}^{\prime\prime}|-\frac{|\lambda_{I}^{\prime}|}{2\pi}|)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\equiv I_{1}+I_{2}.$
Noting that
(4.62) $\qquad\quad\theta(|\lambda_{R}^{\prime}|-C_{0})\theta(||\xi^{\prime\prime}_{n}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-1)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\\ \times|(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})|\geq t/C,\qquad\quad$
we conclude
(4.63) $|I_{1}|\leq o(t^{-1}).$
On the other hand, in view of
$|\lambda_{R}^{\prime}|>C_{0}\gg 1,\quad|\lambda^{\prime}_{I}|<1,\quad b=(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}\sim|\lambda_{R}^{\prime}|,$
we have that stationary points $\pm b$ are not contained in the support of $\theta(1-||\xi^{\prime\prime}_{n}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)$ and
(4.64) $|\partial_{\xi^{\prime\prime}_{n}}\mathfrak{G}|\geq{{\lambda^{\prime}}_{R}^{2}}/C.$
Hence applying integration by parts with respect to $\xi^{\prime\prime}_{n}$, we obtain
(4.65) $|I_{2}|\leq\mathcal{O}(t^{-1}).$
Note that, even we apply argument as in the proof of Proposition 3.1, taking $\xi^{\prime\prime}_{n}$-derivatives on the cut off function $\theta(-(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$ yields a contribution of order $\mathcal{O}(t^{-1})$ rather than $o(t^{-1})$ in (4.65).
[…]
(4.66) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime})\Xi(\lambda^{\prime})[1-\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})]$
$\times$ $\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|\leq o(t^{-1}).$
Without loss of generality, the analysis of (4.66) can be reduced to cases:
  • (1”)
    $\psi_{r/C,0}(\lambda^{\prime}_{R})\neq 0$ and $\psi_{r,r}(\lambda^{\prime}_{I})=0$;
  • (2”)
    $\psi_{r/C,0}(\lambda^{\prime}_{R})=0$.
For Case (1”), $|\partial_{\lambda^{\prime}_{I}}[{\lambda^{\prime}}^{3}_{I}+(a-3{\lambda^{\prime}}_{R}^{2})\lambda^{\prime}_{I}]|>r/C^{\prime}_{0}$ for some positive constant $C_{0}^{\prime}$. Hence the proof of (4.66) can be established by applying integration by parts with respect to $\lambda_{I}^{\prime}$ and $|\lambda_{R}^{\prime}|<r$.
For Case (2”), the proof of (4.66) can be obtained by applying integration by parts with respect to $\lambda^{\prime}$, Lemma 4.5, the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, (4.57), and an adaptation of the argument in the proof of Proposition 3.1.
Lemma 4.9.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(4.67) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\Xi(\lambda^{\prime})\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|\leq o(t^{-8/9}).$
Proof.
Consider the decomposition
(4.68) ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|{\lambda^{\prime}}_{R}|-t^{-5/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})$
$=$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|{\lambda^{\prime}}_{R}|-t^{-5/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-3.5/9})$
$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-1.9/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)$
$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|).$
From $\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})\neq 0$}, we can show that the $L^{1}(d\xi^{\prime\prime}_{n})$-norms of the amplitude function $\mathcal{F}^{(n)}$ over the corresponding domains for the first and second terms on the right-hand side of (4.68) are $o(t^{-1})$. Along with (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, and (2.12), the analysis reduces to studying the contribution over the domain corresponding to the third term, which is bounded by:
(4.69) $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
$+$ $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)(1-\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime}))\mathfrak{C}\mathfrak{T}_{0,(n)}\psi_{t^{-0.95/3},b}(\xi_{n}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
$+$ $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)(1-\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime}))\mathfrak{C}\mathfrak{T}_{0,(n)}(1-\psi_{t^{-0.95/3},b}(\xi_{n}^{\prime\prime}))\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
$\equiv$ $I_{1}+I_{2}+I_{3}.$
Applying Proposition 4.1, and using estimates $|\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime})|_{L^{1}(d\lambda_{I}^{\prime})}\leq Ct^{-0.9/3}$, $|\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)|_{L^{1}(d\lambda_{R}^{\prime})}\leq Ct^{-1.9/9}$, and (2.3), we obtain
(4.70) $|I_{1}|\leq Ct^{-1.9/9-0.9/3-1/2}=o(t^{-1}).$
Moreover, using the two stationary points $\pm b=\pm(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$ of $\mathfrak{S}$, we have
(4.71) $(1-\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime}))\psi_{t^{-0.95/3},b}(\xi_{n}^{\prime\prime})|(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})|\geq\frac{1}{C}t^{-1/3},$
which implies
(4.72) $(1-\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime}))\psi_{t^{-0.95/3},b}(\xi_{n}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $|(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})|\geq t^{1-5/9-3/9}/C.$
Together with (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, and (2.12), yields
(4.73) $|I_{2}|\leq o(t^{-1}).$
Finally, for $I_{3}$, we apply integration by parts with respect to $\xi_{n}^{\prime\prime}$, (4.26), $b=(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$, and $|\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)|_{L^{1}(d\lambda_{R}^{\prime})}\leq Ct^{-1.9/9}$ to conclude
(4.74) $|I_{3}|\leq Ct^{-1+0.95/3-1.9/9}=o(t^{-8/9}).$
The obstruction to the $o(t^{-1})$ estimate arises when $\xi_{n}^{\prime\prime}$ lies in the vicinity of $\pm\frac{\lambda_{I}^{\prime}}{2\pi}$, and when $\lambda^{\prime}_{I}$ in near $\pm b$, stationary points of the phase function $\mathfrak{S}$:
Lemma 4.10.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(4.75) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\widetilde{s}_{c}(\lambda^{\prime})e^{\beta_{n+1}2\pi itS_{0}}(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\Xi(\lambda^{\prime})\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\quad\times\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(t^{-7/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\leq Ct^{-8/9}.$
Proof.
Decompose
(4.76) LHS of (4.75)
$\leq$ $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\psi_{t^{-1/9},b}(\lambda_{I}^{\prime})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(t^{-7/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
$+$ $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $[1-\psi_{t^{-1/9},b}(\lambda_{I}^{\prime})]\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(t^{-7/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\equiv I+I^{\prime}.$
Using (2.3), the factor $\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})$}, and (2.12),
(4.77) $I\leq Ct^{-1/9-7/9}.$
Integration by parts with respect to $\lambda_{I}^{\prime}$, using (2.3), the factor $\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})$}, $b=(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$, and (4.41), we obtain
(4.78) $I^{\prime}\leq Ct^{-1+1/9}.$
As $m(x,\lambda^{\prime})$ depends nontrivially on the entire complex plane and integrands of $\lambda^{\prime}$-derivatives of the Cauchy integrals behave like Dirac functions near $\pm\lambda_{I}^{\prime}/2\pi$. This means that, even though we’re estimating in a narrow region of size $t^{-7/9}$ around these points, the contributions to the $\lambda^{\prime}_{I}$-derivative estimates remain of order $(1+|\lambda^{\prime}_{R}|)\mathcal{O}(1)$ (see (4.78)). Hence, although the estimate $t^{-8/9}$ is not optimal, there are limitations to improving it using our current approach. These limitations arise from constraints in the estimate for $I_{3}$ in (4.74), which requires the two exponents–specifically, $t^{-1.9/9}$ and $t^{-0.95/3}$–to be close to each other. At the same time, the estimate for $I_{1}$ in (4.70) imposes the condition that the product of the two corresponding factors must be close to $t^{-1/2}$.
Therefore, together with Proposition 4.2, Lemmas 4.8 - 4.10, yields:
Theorem 5.
Assume the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,
$u_{2,0}\sim o({t}^{-8/9}).$
5. Long time asymptotics of $u_{2,1}(x)$
We adapt the approach from Section 4 to derive the asymptotic behavior of $u_{2,1}$. To facilitate integration by parts without imposing additional conditions on $\partial^{j}_{\lambda^{\prime}_{I}}\widetilde{s}_{c}$ and $\lambda^{\prime}\widetilde{s}_{c}$ near $\lambda^{\prime}_{I}=0$ (cf [5]), particular care is needed, and the argument becomes more involved.
Throughout this section, $a$, $r$, $_{i}, tt$ are as defined in Definition 1.
5.1. The Cauchy integrals
5.1.1. Representation formulas of the Cauchy integrals
In this subsection, we present two distinct representations of $\partial_{x_{1}}\widetilde{(\mathcal{C}T)^{n}1}$. Each representation is useful in a different context.
Lemma 5.1.
If the assumption of Theorem 1 holds for $u_{0}\in{\mathfrak{M}^{1,q}}$ then
(5.1) $\partial_{x_{1}}\widetilde{\mathcal{C}T1}(t_{1},t_{2},t;\lambda^{\prime})$
$=$ $e^{i\pi tS_{0}(a;\lambda^{\prime})}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ \left(\partial_{x^{\prime}_{1}}[u_{0}\mathfrak{m}_{0}]\right)(x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
$\times$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{1}^{\prime},x_{2}^{\prime};\xi_{1}^{\prime\prime})\equiv e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{1,(1)}1,$
or
(5.2) $\partial_{x_{1}}\widetilde{\mathcal{C}T1}(t_{1},t_{2},t;\lambda^{\prime})$
$=$ $e^{i\pi tS_{0}(a;\lambda^{\prime})}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ \left(\partial_{x^{\prime}_{1}}[u_{0}\mathfrak{m}_{0}]\right)(x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
$\times$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{1}^{\prime},x_{2}^{\prime};\xi_{1}^{\prime\prime})[1-\psi_{1,\frac{\lambda_{I}^{\prime}}{2\pi}}(\xi_{1}^{\prime\prime})]$
$+$ $e^{i\pi tS_{0}(a;\lambda^{\prime})}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ [u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
$\times$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{1}^{\prime},x_{2}^{\prime};\xi_{1}^{\prime\prime})\psi_{1,\frac{\lambda_{I}^{\prime}}{2\pi}}(\xi_{1}^{\prime\prime})\cdot{\color[rgb]{0,0,0}(2\pi i)}(\xi_{1}^{\prime\prime}-\frac{\lambda_{I}^{\prime}}{2\pi})$
$\equiv$ $e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{1,(1)}[1-\psi_{1,\frac{\lambda_{I}^{\prime}}{2\pi}}(\xi_{1}^{\prime\prime})]+e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{0,(1)}\psi_{1,\frac{\lambda_{I}^{\prime}}{2\pi}}(\xi_{1}^{\prime\prime})\cdot{\color[rgb]{0,0,0}(2\pi i)}(\xi_{1}^{\prime\prime}-\frac{\lambda_{I}^{\prime}}{2\pi}),$
with $\mathfrak{m}_{0}$ satisfying (4.5), is holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when $\lambda^{\prime}_{I}\neq 0$.
Moreover,
(5.3) $\partial_{x_{1}}\widetilde{(\mathcal{C}T)^{n}1}(t_{1},t_{2},t;\lambda^{\prime})\equiv$ $e^{\beta_{n}i\pi tS_{0}(a;\lambda^{\prime})}\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
is holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when ${\lambda^{\prime}}_{I}\neq 0$. Here
(5.4) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
$=$ $\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\mathfrak{C}\mathfrak{T}_{1,(h)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
or
(5.5) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
$=$ $\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}$
$\times$ $\{\,\mathfrak{C}\mathfrak{T}_{1,(h)}[1-\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})]+\mathfrak{C}\mathfrak{T}_{0,(h)}\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})\cdot{\color[rgb]{0,0,0}(2\pi i)}(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\,\}$
$\times$ $\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
where $\xi^{\prime\prime}_{n+1}=\frac{\lambda_{I}^{\prime}}{2\pi}$.
Finally,
(5.6) $|\partial_{\lambda^{\prime}_{I}}\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}|\leq C(1+|\lambda_{R}^{\prime}|).$
Proof.
Using the representation formulas (5.1) and (5.4), the proof proceeds by the same argument as in Lemma 4.2 and 4.4.
Note that when $n=1$, (5.5) and (5.4) reduce to (5.2) and (5.1) respectively upon identifying that $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}=\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}=1$ and $\xi^{\prime\prime}_{n+1}=\frac{\lambda_{I}^{\prime}}{2\pi}$. For brevity, we will henceforth use (5.3)-(5.5) to denote $\partial_{x_{1}}\widetilde{(\mathcal{C}T)^{n}1}$ for all $n\geq 1$.
5.1.2. Asymptotics of the Cauchy integrals
Proposition 5.1.
If the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{1,q}$ then for $\color[rgb]{0,0,0}|a|>\delta>0$,
(5.7) $|\partial_{x_{1}}\widetilde{\left(\mathcal{C}T\right)^{n}1}|\leq\mathcal{O}(t^{-1/2}),\quad\textit{ as $t\to\infty$}.$
Proof.
Using the representation formula (5.4), the proof proceeds by the same argument as in Proposition 4.1. ∎
Without the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, it becomes more difficult to justify when the integrability of $(1+|\lambda^{\prime}|)\widetilde{s}_{c}$ holds. To address this, we will require the following reduction results (cf. Proposition 4.2).
Proposition 5.2.
Suppose the assumption of Theorem 1 is valid for $u_{0}\in\mathfrak{M}^{3,q}$, $\Xi(\lambda^{\prime})$ defined by (4.56), and $|a|>\delta>0$. Then, as $t\to\infty$,
(5.8) $|u_{2,1}(x)|\leq$ $C\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(1-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})$
$\times$ $\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|+\mathcal{O}(t^{-1}).$
Proof.
Noting that
(5.9) $\qquad\quad\theta(|\lambda_{R}^{\prime}|-t^{-5/9})[1-\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})]\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\\ \times|(x_{2,h}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})(\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime})|\geq t^{1-5/9}/C,\qquad\quad$
applying Proposition 5.2, the proof follows the argument used in Proposition 4.2 and Step 1 of Lemma 4.8.
5.2. Long time asymptotics of $u_{2,1}(x)$ when $\color[rgb]{0,0,0}a>+\delta>0$
Throughout this subsection, we assume $\color[rgb]{0,0,0}a>+\delta>0$, and define the parameters $\psi_{r,w_{0}}$ and $u_{2,1}$ as in (3.20) and (1.5), respectively. We also set $b=(-r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$ and adopt the terminology established in Lemma 5.1.
Building on Proposition 5.2, we will decompose the estimates for $u_{2,1}$ into two parts, depending on whether $||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||>t^{-6/9}$ or not. Precisely,
Lemma 5.2.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(5.10) $u_{2,1}(x)\leq$ $C\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\sum_{h=1}^{n}(P_{n,h}^{>}+P_{n,h}^{<})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|+\mathcal{O}(t^{-1}),$
where
(5.11) $P_{n,h}^{>}=$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(1-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-6/9}),$
$P_{n,h}^{<}=$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\{\ \mathfrak{C}\mathfrak{T}_{0,(h+1)}(-2\xi_{h+1}^{\prime\prime})$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h+1}^{\prime}|)\theta(1-||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||)\theta(||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||-t^{-6/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(t^{-6/9}-|\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime}|)\theta(|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|-t^{-6/9})\ \}.$
Here, for brevity, when $h=n$, we identify
(5.12) $\qquad\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}(-2\xi_{h+1}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h+1}^{\prime}|)\\ \theta(1-||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||)\theta(||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||-t^{-6/9})=-\frac{\lambda_{I}^{\prime}}{\pi}.\qquad$
Proof.
It reduces to studying
$\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(1-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)$
$\times$ $\theta(t^{-6/9}-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|,$
which is less than
(5.13) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t^{-6/9}-|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$+$ $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|-t^{-6/9})\theta(t^{-6/9}-|\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime}|)(\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime})$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$+$ $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|-t^{-6/9})\theta(t^{-6/9}-|\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime}|)(-2\xi_{h+1}^{\prime\prime})$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$\equiv$ $\sum_{n=1}^{\infty}\sum_{h=1}^{n}Q_{n,h}^{>,-}+\sum_{n=1}^{\infty}\sum_{h=1}^{n}Q_{n,h}^{>,+}+\sum_{n=1}^{\infty}\sum_{h=1}^{n}Q_{n,h}^{<}.$
Using (4.57) and $|(\xi_{h}^{\prime\prime}\pm\xi_{h+1}^{\prime\prime})\theta(t^{-6/9}-|\xi_{h}^{\prime\prime}\pm\xi_{h+1}^{\prime\prime}|)|_{L^{1}(d\xi_{h}^{\prime\prime})}\leq C(t^{-6/9\times 2})$, we obtain
(5.14) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}Q_{n,h}^{>,\pm}\leq Ct^{-6/9\times 2}.$
Applying the above argument, we have
(5.15) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}Q_{n,h}^{<}$
$\leq$ $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}(-2\xi_{h+1}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h+1}^{\prime}|)$
$\times$ $\theta(1-||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||)\theta(||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||-t^{-6/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(t^{-6/9}-|\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime}|)\theta(|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|-t^{-6/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$+$ $Ct^{-6/9\times 2}+o(t^{-1})$
$=$ $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})P_{n,h}^{<}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|+o(t^{-1}).$
The next lemma allows us to restrict our attention to the regime $|\lambda_{R}^{\prime}|>r/C$, which is a weaker condition than requiring $\lambda^{\prime}$ to lie in the support of $\chi(\lambda^{\prime})$ (cf. Lemma 4.6). Nevertheless, it is sufficient for deriving asymptotics away from the vicinity of $\pm\xi^{\prime\prime}_{h+1}$.
Lemma 5.3.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(5.16) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9}){\color[rgb]{0,0,0}\theta(a-3{\lambda^{\prime}_{R}}^{2})}$
$\times$ $\sum_{h=1}^{n}(P_{n,h}^{>}+P_{n,h}^{<})[1-\psi_{r,r}(\lambda^{\prime}_{R})\psi_{5r,b}(2\pi\xi_{h+1}^{\prime\prime})]\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-1}).$
Proof.
By assumption there is no stationary point and $|{\lambda_{R}^{\prime}}|\leq r$, and the analysis can be reduced to cases:
  • (1+)
    $\psi_{r,r}(\lambda^{\prime}_{R})\neq 0$ and $\psi_{5r,b}(2\pi\xi_{h+1}^{\prime\prime})=0$;
  • (2+)
    $\psi_{r,r}(\lambda^{\prime}_{R})=0$.
Notice that $\partial_{\xi^{\prime\prime}_{h+1}}\mathfrak{S}(a;\lambda_{R}^{\prime};2\pi\xi_{h+1}^{\prime\prime})=+12\pi^{2}(2\pi\xi_{h+1}^{\prime\prime})^{2}+(a-3{\lambda^{\prime}}_{R}^{2})\geq r/C$ for both cases. Therefore, integration by parts with respect to $\xi_{h+1}^{\prime\prime}$, using $|{\lambda_{R}^{\prime}}|\leq r$, Lemmas 3.3 and 4.5 (cf. Lemma 4.6), we prove the lemma.
Lemma 5.4.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(5.17) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(a-3{\lambda^{\prime}_{R}}^{2})}$
$\times$ $\sum_{h=1}^{n}P_{n,h}^{>}\psi_{r,r}(\lambda^{\prime}_{R})\psi_{5r,b}(2\pi\xi_{h+1}^{\prime\prime})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-1}),$
(5.18) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}$
$\times$ $\sum_{h=1}^{n}P_{n,h}^{>}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-1}).$
Proof.
We will first discard terms with rapidly decaying amplitudes. Then, through a refined decomposition, we derive the necessary estimates by leveraging the smallness of the integration domains and the factor or $(\xi_{h}^{\prime\prime}\pm\xi_{h+1}^{\prime\prime})$. Integration by parts is not required in the proof.
To prove (5.18), decompose
(5.19) $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-6/9})$
$=$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-4.4/9})$
$+$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(t^{-4.4/9}-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-6/9}).$
Thanks to $\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))$}, $|\lambda_{R}^{\prime}|>r/2$ as $t\gg 1$. Hence the $L^{1}(d\xi^{\prime\prime}_{h})$-norm of the amplitude function $\mathcal{F}^{(h)}$ on the corresponding domain of the first term is less than $o(t^{-1})$. Together with (4.57) and Lemma 5.2, it reduces to showing
(5.20) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}$
$\times$ $P_{n,h}^{>}\theta(t^{-4.4/9}-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}\sim o(t^{-1}).$
Notice
(5.21) LHS of (5.20)
$\leq$ $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\psi_{t^{-2.5/9},0}(2\pi\xi_{h+1}^{\prime\prime})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t^{-4.4/9}-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)$
$\times$ $\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-6/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$+$ $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}[1-\psi_{t^{-2.5/9},0}(2\pi\xi_{h+1}^{\prime\prime})]$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t^{-4.4/9}-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)$
$\times$ $\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-6/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\equiv II_{1}+II_{2}.$
Using
$|(\psi_{t^{-2.5/9},0}(2\pi\xi_{h+1}^{\prime\prime})|_{L^{1}(d\xi_{h+1}^{\prime\prime})}\leq Ct^{-2.5/9},$
(5.22) $|(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\theta(t^{-4.4/9}-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)|_{L^{1}(d\xi_{h}^{\prime\prime})}\leq C(t^{-4.4/9}+|\xi_{h+1}^{\prime\prime}|)t^{-4.4/9},$
and (4.57), we obtain
(5.23) $\begin{split}|II_{1}|\leq C\left(\mathcal{O}(t^{-2.5/9\times 2-4.4/9})+\mathcal{O}(t^{-2.5/9-4.4/9\times 2})\right).\end{split}$
Besides, on the support of $(1-\psi_{t^{-2.5/9},0}(2\pi\xi_{h+1}^{\prime\prime}))$, distance between $\pm\xi_{h+1}^{\prime\prime}$ is greater than $\mathcal{O}(t^{-2.5/9})$. Combining with $|\lambda_{R}^{\prime}|>r$ on the support of $\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))$},
$(1-\psi_{t^{-2.5/9},0}(2\pi\xi_{h+1}^{\prime\prime}))\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|){\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}\\ \times\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-6/9})|(x_{2,h}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})(\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime})|\geq t^{1-6/9-2.5/9}/C,$
which, together with (4.57), implies
(5.24) $\begin{split}|II_{2}|\leq o(t^{-1}).\end{split}$
Therefore, (5.20) is justified.
Since $|\lambda_{R}^{\prime}|>r/C$ is assured by the factor $\psi_{r,r}(\lambda^{\prime}_{R})$. We can prove (5.17) by analogy.
The following lemma shows that the obstruction to obtaining an $o(t^{-1})$ estimate for $u_{2,1}$ lies in the vicinity of $\xi_{h+1}^{\prime\prime}$.
Lemma 5.5.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(5.25) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(a-3{\lambda^{\prime}_{R}}^{2})}$
$\times$ $P_{n,h}^{<}\psi_{r,r}(\lambda^{\prime}_{R})\psi_{5r,b}(2\pi\xi_{h+1}^{\prime\prime})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})},$
(5.26) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}$
$\times$ $P_{n,h}^{<}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.$
Proof.
From (4.57) and (5.19), to prove (5.26), it reduces to justifying
(5.27) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(-(a-3{\lambda^{\prime}_{R}}^{2}))}$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}(-2\xi_{h+1}^{\prime\prime})\theta(1-||\xi_{h+2}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\theta(t|\lambda_{R}^{\prime}|-|x_{2,h+1}^{\prime}|)$
$\times$ $\theta(||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||-t^{-6/9})\theta(t^{-4.4/9}-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(t^{-6/9}-|\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime}|)\theta(|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|-t^{-6/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$\leq$ ${\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.$
To this aim, decomposing $2\xi_{h+1}^{\prime\prime}=-2(\xi_{h+1}^{\prime\prime}-\xi_{h+2}^{\prime\prime})+2\xi_{h+2}^{\prime\prime}$, applying Lemma 5.4, an induction, and Theorem 4, we have
[…]
In an entirely similar way, we can prove (5.25). ∎
Combining Proposition 5.2, Lemmas 5.2-5.5, we conclude:
Theorem 6.
Assume the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,
(5.29) $u_{2,1}\sim{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.$
5.3. Long time asymptotics of $u_{2,1}(x)$ when $\color[rgb]{0,0,0}a<-\delta<0$
Throughout this section, we assume the hypotheses of Theorem 1, $\color[rgb]{0,0,0}a<-\delta<0$, and define the parameters $a$, $r$, $_{i}, tt$, $\psi_{r,w_{0}}$ as in (2.11), (2.15), (2.17), and (3.20) respectively. We also set $b=(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$ and adopt the terminology established in Lemma 5.1.
Similarly, building on Proposition 5.2, we can decompose the estimates for $u_{2,1}$ into two parts, depending on whether $||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||>t^{-7/9}$ or not. Precisely,
Lemma 5.6.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
(5.30) $u_{2,1}(x)\leq$ $C\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $\sum_{h=1}^{n}(\mathbb{P}_{n,h}^{>}+\mathbb{P}_{n,h}^{<})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|+\mathcal{O}(t^{-1}),$
where
(5.31) $\mathbb{P}_{n,h}^{>}=$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(1-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9}),$
$\mathbb{P}_{n,h}^{<}=$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\{\ \mathfrak{C}\mathfrak{T}_{0,(h+1)}(-2\xi_{h+1}^{\prime\prime})$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h+1}^{\prime}|)\theta(1-||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||)\theta(||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||-t^{-7/9})$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(t^{-7/9}-|\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime}|)\theta(|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|-t^{-7/9})\ \}.$
Here, for brevity, when $h=n$, we identify
(5.32) $\qquad\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}(-2\xi_{h+1}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h+1}^{\prime}|)\\ \times\theta(1-||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||)\theta(||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||-t^{-7/9})=-\frac{\lambda_{I}^{\prime}}{\pi}.\qquad$
Proof.
The proof proceeds by the same argument as in Lemma 5.2. ∎
Analogous to Lemma 4.9, we have:
Lemma 5.7.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
$\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathbb{P}_{n,h}^{>}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-8/9}).$
Proof.
Applying Proposition 5.2 and the argument used in Step 2 of Lemma 4.8, it reduces to showing
(5.33) $\color[rgb]{0,0,0}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathbb{P}_{n,h}^{>,\sharp}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-8/9}),$}
where
(5.34) ${\color[rgb]{0,0,0}\mathbb{P}_{n,h}^{>,\sharp}}=$ ${\color[rgb]{0,0,0}\psi_{r/C,0}({\lambda^{\prime}}_{R})\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\psi_{b,b}(\xi_{h+1}^{\prime\prime})\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(1-||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})}$
$\times$ $\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9}).$
To prove the lemma, we follow the approach used in the proof of Lemma 4.9. As in that case, we begin by refining the decomposition before applying integration by parts.
Consider the decomposition
(5.35) ${\color[rgb]{0,0,0}\psi_{b,b}(\xi_{h+1}^{\prime\prime})}\theta(|{\lambda^{\prime}}_{R}|-t^{-5/9})\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9})$
$=$ ${\color[rgb]{0,0,0}\psi_{b,b}(\xi_{h+1}^{\prime\prime})}\theta(|{\lambda^{\prime}}_{R}|-t^{-5/9})\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-3.5/9})$
$+$ ${\color[rgb]{0,0,0}\psi_{b,b}(\xi_{h+1}^{\prime\prime})}\theta(|\lambda_{R}^{\prime}|-t^{-1.9/9})\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)$
$+$ ${\color[rgb]{0,0,0}\psi_{b,b}(\xi_{h+1}^{\prime\prime})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)$
$\times$ $\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||).$
From $b=(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$, we can prove the $L^{1}(d\xi_{h}^{\prime\prime})$-norm of $\mathcal{F}^{(h)}$ on the corresponding domains for the first and the second terms on the right hand side of (5.35) are less than $o(t^{-1})$. Together with (4.57), it then reduces to proving:
(5.36) $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)$
$\times$ $\mathbb{P}_{n,h}^{>,{\color[rgb]{0,0,0}\sharp}}\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-8/9}).$
Notice
LHS of (5.36)
$\leq$ $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(|\lambda^{\prime}_{I}|-1)}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\psi_{t^{-0.9/3},b}(\xi_{h+1}^{\prime\prime})\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)$
$\times$ $\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$+$ $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(|\lambda^{\prime}_{I}|-1)}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}{\color[rgb]{0,0,0}\psi_{b,b}(\xi_{h+1}^{\prime\prime})}(1-\psi_{t^{-0.9/3},b}(\xi_{h+1}^{\prime\prime}))\mathfrak{C}\mathfrak{T}_{0,(h)}\psi_{t^{-0.95/3},b}(\xi_{h}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)$
$\times$ $\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$+$ $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(|\lambda^{\prime}_{I}|-1)}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}{\color[rgb]{0,0,0}\psi_{b,b}(\xi_{h+1}^{\prime\prime})}(1-\psi_{t^{-0.9/3},b}(\xi_{h+1}^{\prime\prime}))\mathfrak{C}\mathfrak{T}_{0,(n)}(1-\psi_{t^{-0.95/3},b}(\xi_{h}^{\prime\prime}))\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)$
$\times$ $\theta(||\xi_{h}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-|\xi_{h+1}^{\prime\prime}||)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$\equiv$ $I_{1}+I_{2}+I_{3}.$
Using $|\psi_{t^{-0.9/3},b}(\xi_{h+1}^{\prime\prime})|_{L^{1}(d\xi_{h+1}^{\prime\prime})}\leq Ct^{-0.9/3}$, $|\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)|_{L^{1}(d\lambda_{R}^{\prime})}\leq Ct^{-1.9/9}$, and Proposition 5.1,
(5.37) $|I_{1}|\leq Ct^{-1.9/9-0.9/3-1/2}=o(t^{-1}).$
Moreover, using the two stationary points $\pm b=\pm\frac{\sqrt{{\lambda^{\prime}}_{R}^{2}+r^{2}}}{2\pi}\geq r$ of $\mathfrak{S}$, we have
(5.38) $(1-\psi_{t^{-0.9/3},b}(\xi_{h+1}^{\prime\prime}))\psi_{t^{-0.95/3},b}(\xi_{h}^{\prime\prime})|(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})(\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime})|\geq\frac{1}{C}t^{-1/3},$
and, then
(5.39) $(1-\psi_{t^{-0.9/3},b}(\xi_{h+1}^{\prime\prime}))\psi_{t^{-0.95/3},b}(\xi_{h}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
$\times$ $|(x_{2,h}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})(\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime})|\geq t^{1-5/9-1/3}/C.$
Consequently, the $L^{1}(d\xi_{h}^{\prime\prime})$}-norm of the amplitude function $\mathcal{F}^{(h)}$ on the corresponding domain is less than $o(t^{-1})$, and
(5.40) $|I_{2}|\leq o(t^{-1}).$
Finally, for $I_{3}$, integration by parts with respect to $\xi_{h}^{\prime\prime}$, using (5.6), $|\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)|_{L^{1}(d\lambda_{R}^{\prime})}\leq Ct^{-1.9/9}$, and $b=(r^{2}+{\lambda^{\prime}}_{R}^{2})^{1/2}/{2\pi}$,
(5.41) $|I_{3}|\leq\mathcal{O}(t^{-1+0.95/3-1.9/9})\leq o(t^{-8/9}).$
Consequently, (5.36) follows from (5.37), (5.40), and (5.41).
Analogous to Lemma 4.10, we have:
Lemma 5.8.
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
$\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathbb{P}_{n,h}^{<}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-8/9}).$
Proof.
From (4.57), (5.35) and a similar argument as argument used in Step 2 of Lemma 4.8, to prove the lemma, it reduces to justifying
(5.42) $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}(-2\xi_{h+1}^{\prime\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda_{R}^{\prime})\psi_{b,b}(\xi_{h+1}^{\prime\prime})}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h+1}^{\prime}|)$
$\times$ $\theta(||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{h+1}^{\prime\prime}|-|\xi_{h+2}^{\prime\prime}||)$
$\times$ $\mathfrak{C}\mathfrak{T}_{0,(h)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,h}^{\prime}|)\theta(t^{-7/9}-|\xi_{h}^{\prime\prime}+\xi_{h+1}^{\prime\prime}|)\theta(|\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|$
$\leq$ $o(t^{-8/9}).$
To this aim, via decomposing $2\xi_{h+1}^{\prime\prime}=-2(\xi_{h+1}^{\prime\prime}-\xi_{h+2}^{\prime\prime})+2\xi_{h+2}^{\prime\prime}$, applying Lemma 5.7, an induction, and Theorem 5, we have
[…]
Thanks to Theorem 5 and the preceding remark, we observe that the estimate $t^{-8/9}$ is essentially optimal within the context of this approach. Therefore, combining Lemmas 5.6-5.8, we obtain:
Theorem 7.
Assume the assumption of Theorem 1 holds for $u\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,
$u_{2,1}\sim o(t^{-8/9}).$
Appendix A A technical lemma
We provide one key estimate used in the derivation of new representation formulas.
Lemma A.1.
Suppose (2.1) is true. Let $\mathfrak{m}_{0}(x_{1},x_{2})$ be defined by (4.5). For $j=0,1$,
(A.1) $|\partial_{x_{1}}^{j}(\mathfrak{m}_{0}-1)|_{L^{\infty}}\leq|\left(\partial_{x_{1}}^{j}(\mathfrak{m}_{0}(x_{1},x_{2};\overline{\zeta(\xi)})-1)\right)^{\wedge_{x_{1},x_{2}}}|_{L^{1}(d\xi_{1}d\xi_{2})}\leq C.$
Proof.
We will adpt the proof given in [8]. From (2.7), for $j=0,1$,
(A.2) $\begin{split}\left[\partial_{x_{1}}^{j}(m_{0}(x_{1},x_{2};\lambda)-1)\right]^{\wedge_{x_{1},x_{2}}}(\xi;\lambda)=&\left[\mathcal{C}T(2\pi i\xi_{1})^{j}(m_{0}(x_{1},x_{2};\lambda)-1)\right]^{\wedge_{x_{1},x_{2}}}(\xi;\lambda)\\ &+\left[\mathcal{C}T(2\pi i\xi_{1})^{j}\right]^{\wedge_{x_{1},x_{2}}}(\xi;\lambda).\end{split}$
Applying the Fourier theory and (4.3) and Theorem 2, we obtain
(A.3) $\begin{split}|\left[\mathcal{C}T(2\pi i\xi_{1})^{j}\right]^{\wedge_{x_{1},x_{2}}}(\xi;\lambda)|_{L^{1}(d\xi_{1}d\xi_{2})}=&|\frac{(2\pi i\xi_{1})^{j}s_{c}}{p_{\lambda}(\xi)}|_{L^{1}(d\xi_{1}d\xi_{2})}\leq C|\xi_{1}^{j}s_{c}|_{L^{\infty}\cap L^{2}(d\xi_{1}d\xi_{2})}\\ \leq&C\sum_{|l|\leq 2+j}|\partial_{x}^{l}u_{0}|_{L^{1}\cap L^{2}},\end{split}$
and
(A.4) $\left[\mathcal{C}T(2\pi i\xi_{1})^{j}f\right]^{\wedge_{x_{1},x_{2}}}(\xi_{0};\lambda)$
$=$ $\iint\left[\frac{1}{2\pi i}\iint\frac{(2\pi i\xi_{1})^{j}s_{c}(\zeta)f(x_{1},x_{2};\overline{\zeta})e^{2\pi i(x_{1}\xi_{0,1}+x_{2}\xi_{0,2})}}{\lambda-\zeta}d\overline{\zeta}\wedge d\zeta\right]dx_{1}dx_{2}$
$=$ $\frac{1}{2\pi i}\iint\frac{(2\pi i\xi_{1})^{j}s_{c}(\zeta)}{\lambda-\zeta}\widehat{f}(\xi_{1}-\xi_{0,1},\xi_{2}-\xi_{0,2};\overline{\zeta})d\overline{\zeta}\wedge d\zeta\equiv\mathbf{R}_{(2\pi i\xi_{1})^{j}s_{c}}\widehat{f}(\xi_{0};\lambda).$
In view of (4.3), Theorem 2, and the Minkowski inequality,
(A.5) $|\mathbf{R}_{(2\pi i\xi_{1})^{j}s_{c}}\widehat{f}(\xi_{0};\lambda)|_{L^{1}(d\xi_{0,1}d\xi_{0,2})}\leq C|\widehat{f}|_{L^{1}(d\xi_{1}d\xi_{2})}.$
Combining (A.2)-(A.5), and the Minkowski inequality, we obtain
(A.6) $|[\partial_{x_{1}}^{j}(m_{0}(x_{1},x_{2};\lambda)-1)]^{\wedge_{x_{1},x_{2}}}(\xi;\lambda)|_{L^{1}(d\xi_{1}d\xi_{2})}\leq C|\frac{\xi_{1}^{j}s_{c}}{p_{\lambda}}|_{L^{1}(d\xi_{1}d\xi_{2})}\leq C\sum_{|l|\leq 2+j}|\partial_{x}^{l}u_{0}|_{L^{1}\cap L^{2}}.$
Using the definition of Riemann sums,
$|[\partial_{x_{1}}^{j}(m_{0}(x_{1},x_{2};\overline{\zeta(\xi)})-1)]^{\wedge_{x_{1},x_{2}}}|_{L^{1}(d\xi_{1}d\xi_{2})}$
$\leq$ $\sup_{\lambda}|[\partial_{x_{1}}^{j}(m_{0}(x_{1},x_{2};\lambda)-1)]^{\wedge_{x_{1},x_{2}}}(\xi;\lambda)|_{L^{1}(d\xi_{1}d\xi_{2})}.$
Therefore, (A.1) is justified.
Appendix B List of Symbols
Table B.1. List of Symbols
Notation and Definition Page Notation and Definition Page
Coordinates $\qquad\widetilde{f}(\zeta^{\prime})$, 5
$\qquad x=(x_{1},x_{2},x_{3})$, 4 $\qquad\mathbb{S}_{0}(t_{1},t_{2},\zeta),\ S_{0}(a;\zeta^{\prime})$, 5
$\qquad\partial_{x}^{l}=\partial_{x_{1}}^{l_{1}}\partial_{x_{2}}^{l_{2}}\partial_{x_{3}}^{l_{3}},\ |l|=l_{1}+l_{2}+l_{3}$, 4 $\qquad\nabla S_{0}(a;\zeta^{\prime}),\ \Delta S_{0}(a;\zeta^{\prime})$, 9
$\qquad\xi=(\xi_{1},\xi_{2})$, 4 $\qquad a$, 5
$\qquad C,\ \epsilon_{0},\ \delta$ 4,2 $\qquad\pm r$ stationary point for $S_{0}(\zeta^{\prime})$    6
Potentials (KPII solutions) Special functions
$\qquad u(x),\ u_{0}(x_{1},x_{2})$, 2   Airy function $Ai(z)$, 11
$\qquad u_{1}(x),\ u_{1,1}(x),\ u_{1,2}(x)$, 2,8   Heaviside function $\theta(s)$, 7
$\qquad u_{2,0}(x),\ u_{2,1}(x)$ 2 $\qquad\mathfrak{M}^{p,q}$, 4
$\qquad\psi_{r,w_{0}}(s),\ \chi(\lambda^{\prime})$, 8
Inverse scattering theory $\qquad\Xi(\lambda^{\prime})$ 23
$\qquad\mathcal{S},\ s_{c},\ \mathcal{C},\ T$ 4,5
CIO (new representation)
Fourier transform $\qquad\mathfrak{m}_{0}(x_{1}^{\prime},x_{2}^{\prime}),\ x_{1,n}^{\prime},\ x_{2,n}^{\prime}$, 13,16
$\qquad\widehat{f}(\xi)$, 4 $\qquad\xi_{1}^{\prime\prime},\ \xi_{n}^{\prime\prime},\ \xi_{h}^{\prime\prime},\ \xi_{n+1}^{\prime\prime}$, 13,16,28
$\qquad\phi^{\wedge_{\zeta_{R}^{\prime}}}(\zeta_{R}^{\prime}),\ \phi^{\wedge_{\zeta_{I}^{\prime}}}(\zeta_{I}^{\prime})$, 10 $\qquad[\mathfrak{C}\mathfrak{T}]^{0,(n)},[\mathfrak{C}\mathfrak{T}]^{1,(n)},\mathfrak{C}\mathfrak{T}_{0,(n)},\mathfrak{C}\mathfrak{T}_{1,(n)}$, 13,16,27,28
$\qquad\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime}),\,\mathfrak{S}^{\sharp}(a,t;x_{1}^{\prime},x_{2}^{\prime};\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})$, 13
Stationary theory $\qquad\mathcal{F}(t;\lambda^{\prime};x_{2}^{\prime};\xi_{1}^{\prime\prime}),\ \mathcal{F}^{(n)}(t;\lambda^{\prime};x_{2,n}^{\prime};\xi_{n}^{\prime\prime})$, 13,16
$\qquad(t_{1},t_{2},t)$, 5 $\qquad\beta_{n}$, 16
$\qquad\zeta=\zeta_{R}+i\zeta_{I},\ \zeta^{\prime}=\zeta^{\prime}_{R}+i\zeta^{\prime}_{I}$, 5 $\qquad\pm b$ stationary points for $\mathfrak{G}(\xi_{n}^{\prime\prime})$, 18
$\qquad(\xi_{1}^{\prime},\xi_{2}^{\prime}),\ \partial_{\zeta_{R}^{\prime}},\ \partial_{\zeta_{I}^{\prime}}$, 5 $\qquad P_{n,h}^{>},\ P_{n,h}^{<},\ \mathbb{P}_{n,h}^{>},\ \mathbb{P}_{n,h}^{<}$ 29,34

### v4

@@
 where
-(3.2) $\begin{gathered}G_{u_{0}}f=1+\iint e^{2\pi i(x_{1}\xi_{1}+x_{2}\xi_{2})}\frac{\left[u_{0}f\right]^{\wedge}(\xi_{1},\xi_{2},\lambda)}{p_{\lambda}(\xi_{1},\xi_{2})}d\xi_{1}d\xi_{2},\\ p_{\lambda}(\xi_{1},\xi_{2})=(2\pi i\xi_{1}+\lambda)^{2}-(2\pi i\xi_{2}+{\lambda}^{2}).\end{gathered}$
-Lemma 3.1.
-(3.3) $|G_{u_{0}}f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}|f|_{L^{\infty}},$
-(3.4) $|\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}},\ |\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}},$
-and, for $j\geq 2$,
-(3.5) $|\left[\partial^{j}_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|\lambda_{I}|)^{j-1}(1+|x_{2}|)^{j-1}|u_{0}|_{\mathfrak{M}^{j-1,0}}|f|_{L^{\infty}}+C|\partial^{j-1}_{\lambda_{R}}\left(\lambda_{I}G_{u_{0}}f\right)|,$
-(3.6) $|\left[\partial^{j}_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|\lambda|)^{j-1}(1+|x_{2}|)^{j-1}|u_{0}|_{\mathfrak{M}^{j-1,0}}|f|_{L^{\infty}}+C|\partial^{j-1}_{\lambda_{I}}\left(\lambda_{I}G_{u_{0}}f\right)|,$
-(3.7) $|\left[\partial_{\lambda_{R}}\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|{\color[rgb]{0,0,0}\lambda_{R}}|)(1+|x_{2}|)|u_{0}|_{\mathfrak{M}^{1,0}}|f|_{L^{\infty}}+C|\partial_{\lambda_{I}}\left(\lambda_{I}G_{u_{0}}f\right)|.$
-Proof.
-We can prove (3.3) by applying the estimates [8]:
+(3.2) $\begin{gathered}G_{u_{0}}f=\iint e^{2\pi i(x_{1}\xi_{1}+x_{2}\xi_{2})}\frac{\left[u_{0}f\right]^{\wedge}(\xi_{1},\xi_{2},\lambda)}{p_{\lambda}(\xi_{1},\xi_{2})}d\xi_{1}d\xi_{2},\\ p_{\lambda}(\xi_{1},\xi_{2})=(2\pi i\xi_{1}+\lambda)^{2}-(2\pi i\xi_{2}+{\lambda}^{2}).\end{gathered}$
+Applying the estimates [8]:
 (3.3) $\begin{gathered}\left|\frac{1}{p_{\lambda}}\right|_{L^{1}(\Omega_{\lambda},d\xi_{1}d\xi_{2})}\leq\frac{C}{(1+|\lambda_{I}|^{2})^{1/2}},\quad\left|\frac{1}{p_{\lambda}}\right|_{L^{2}(\Omega_{\lambda}^{c},d\xi_{1}d\xi_{2})}\leq\frac{C}{(1+|\lambda_{I}|^{2})^{1/4}},\end{gathered}$
-where $\Omega_{\lambda}=\{(\xi_{1},\xi_{2})\in{\mathbb{R}}^{2}:|p_{\lambda}(\xi_{1},\xi_{2})|<1\}$.
-For $(x_{2}-x_{2}^{\prime})\neq 0$, via the Fourier inversion formula, the convolution, and the residue theorem, the operator $G$ can be written as [1]:
+where $\Omega_{\lambda}=\{(\xi_{1},\xi_{2})\in{\mathbb{R}}^{2}:|p_{\lambda}(\xi_{1},\xi_{2})|<1\}$, we have
+(3.4) $|G_{u_{0}}f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}|f|_{L^{\infty}}.$
+Moreover, via the Fourier theory and the residue theorem, the operator $G_{u_{0}}$ can be written as [1]:
 (3.5) $G_{u_{0}}f(x_{1},x_{2},\lambda)$
@@
 (3.6) $\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
-$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}\left(4\pi i(x_{2}-x_{2}^{\prime})\xi_{1}\right)$
-$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
 $=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
-$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}{\color[rgb]{0,0,0}+}\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi\lambda_{I})G_{u_{0}}f.$
-Along with (3.3), yields
-$|\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}}.$
-Similarly,
-(3.11) $\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
+$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left[4\pi i\xi_{1}(x_{2}-x_{2}^{\prime})\right]e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
+$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
+$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
+${\color[rgb]{0,0,0}+}$ $\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi\lambda_{I})G_{u_{0}}(x_{2}-x_{2}^{\prime})f.$
+As a result,
+(3.7) $\begin{gathered}\partial_{\lambda_{R}}m_{0}=\sum_{k=0}^{1}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m_{1,k,k^{\prime}},\quad|m_{1,k,k^{\prime}}|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}},\end{gathered}$
+and, for $\lambda^{\prime}_{I}\neq 0$,
+(3.8) $|\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|=|\partial_{\lambda_{R}}s_{c}|$
+$\leq$ $C|\lambda_{I}\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|+C|\left[{u_{0}m_{1,0,0}}\right]^{\wedge}|+C|\lambda_{I}\left[{u_{0}m_{1,1,0}}\right]^{\wedge}|+C|\lambda_{I}\left[{x_{2}u_{0}m_{1,1,1}}\right]^{\wedge}|$
+$\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{1,0}})(1+|\lambda^{\prime}_{I}|).$
+Lemma 3.1.
+Suppose $|t_{2}|<C$ and $\lambda^{\prime}_{I}\neq 0$.
+(3.9) $|\partial_{\lambda^{\prime}_{R}}^{j}\widetilde{s}_{c}|\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{j,0}})(1+|\lambda^{\prime}_{I}|)^{j},$
+(3.10) $|\partial_{\lambda^{\prime}_{I}}^{j}\widetilde{s}_{c}|\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{j,0}})(1+|\lambda^{\prime}|)^{j},$
+(3.11) $|\partial_{\lambda^{\prime}_{R}}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{2,0}})(1+|\lambda^{\prime}|)^{2}.$
+Proof.
+Similarly to the argument for (3.8), for $\lambda^{\prime}_{I}\neq 0$, using
+(3.12) $\left[\partial_{\lambda_{R}}^{2}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
+$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
+$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}\partial^{2}_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
+- $\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
+$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}(4\pi\lambda_{I})(x_{2}-x_{2}^{\prime})\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
+${\color[rgb]{0,0,0}+}$ $\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi i\lambda_{I})\partial_{\lambda_{R}}\left(G_{u_{0}}(x_{2}-x_{2}^{\prime})f\right),$
+$\vdots$
+(3.13) $\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
+$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
 $\times$ $\left[\partial_{\lambda_{I}}\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\right]e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
@@
 $\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\cdot[-4\pi\xi_{1}(x_{2}-x_{2}^{\prime})]e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})},$
+$\vdots$
+we have
+(3.14) $\partial_{\lambda_{R}}\partial_{\lambda_{I}}m_{0}=$ $\sum_{k=0}^{2}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m_{k,k^{\prime}},\quad|m_{k,k^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}}),$
+(3.15) $\partial^{j}_{\lambda_{R}}m_{0}=$ $\sum_{k=0}^{j}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m_{j,k,k^{\prime}},\quad|m_{j,k,k^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}}),$
+(3.16) $\partial^{j}_{\lambda_{I}}m_{0}=$ ${\sum_{k=0}^{j}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m^{+}_{j,k,k^{\prime}}}+\sum_{l=1}^{j-1}\sum_{h=0}^{j-l}\sum_{h^{\prime}=0}^{h}(x_{1}+\lambda_{R}x_{2})^{l}\lambda_{I}^{h}x^{h^{\prime}}_{2}m^{-}_{j,l,h,h^{\prime}},$
+${|m^{+}_{j,k,k^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}}),\quad|m^{-}_{j,l,h,h^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{l+h-h^{\prime},0}}).}$
+Hence the proof of the lemma can be justified by taking derivatives of (2.2).
+We have sharper estimates for the following first derivatives:
+Lemma 3.2.
+Suppose the assumption of Theorem 1 holds. For $\lambda^{\prime}_{I}\neq 0$,
+(3.17) $|\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|\leq C{\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}}),\quad|\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\}){\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}}),$
+(3.18) $|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|\leq C{\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}}),\quad|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\}){\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}}).$
+Proof.
+For $j=1,2$, via (2.4), (3.1), the Fourier theory, and integration by parts,
+(3.19) $|\left[\partial_{x_{j}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C(|u_{0}|_{\mathfrak{M}^{0,1}}|f|_{L^{\infty}}+|u_{0}|_{\mathfrak{M}^{0,0}}|\partial_{x_{j}}f|_{L^{\infty}}),$
+(3.20) $|\partial_{x_{j}}m_{0}|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,1}},$
+(3.21) $|\left[\partial_{x_{j}}G_{u_{0}}\right]m_{0}|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,1}},$
+(3.22) $|\partial_{x_{j}}m_{1,1,k}|_{L^{\infty}}\leq C(1+|u_{0}|_{\mathfrak{M}^{1,1}}),\quad k=0,1.$
+Combining with (3.6) and integration by parts, for $\lambda^{\prime}_{I}\neq 0$, we obtain
+(3.23) $|\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|=|\partial_{\lambda_{R}}s_{c}|$
+$\leq$ $C(|\lambda_{I}\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|+|\left[{u_{0}m_{1,0,0}}\right]^{\wedge}|+|\lambda_{I}\left[{u_{0}m_{1,1,0}}\right]^{\wedge}|+|\lambda_{I}\left[{x_{2}u_{0}m_{1,1,1}}\right]^{\wedge}|)$
+$\leq$ $C(|\left[\partial_{x_{1}}\{x_{2}{u_{0}m_{0}}\}\right]^{\wedge}|+|u_{0}|_{\mathfrak{M}^{0,0}}+|\left[\partial_{x_{1}}\{{u_{0}m_{1,1,0}}\}\right]^{\wedge}|+|\left[\partial_{x_{1}}\{{x_{2}u_{0}m_{1,1,1}}\}\right]^{\wedge}|)$
+$\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}),$
+and, similarly,
+(3.24) $|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|\leq C|\lambda_{I}\partial_{\lambda_{R}}s_{c}|$
+$\leq$ $C|\partial_{\lambda_{R}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{R}}\left[u_{0}(\partial_{x_{1}}m_{0})\right]^{\wedge}|$
+$\leq$ $C|\partial_{\lambda_{R}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{R}}\left[u_{0}(1-G_{u_{0}})^{-1}(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|$
+$\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}).$
+In an entirely similar way, we can justify the corresponding estimates for $\partial_{\lambda_{I}^{\prime}}\widetilde{s}_{c}$ and $\lambda_{I}^{\prime}\partial_{\lambda_{I}^{\prime}}\widetilde{s}_{c}$:
+(3.25) $|\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|=|\partial_{\lambda_{I}}s_{c}|$
+$\leq$ $C|\left[x_{1}{u_{0}m_{0}}\right]^{\wedge}|+C|\lambda_{R}\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|$
+[…]
+$\leq$ $C|[x_{1}{u_{0}m_{0}}]^{\wedge}|+C\min\{\ |\lambda_{R}|\times|\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|,\ \frac{1}{|\lambda_{I}|}\times|\left[\partial_{x_{2}}\{x_{2}{u_{0}m_{0}}\}\right]^{\wedge}|\ \}$
+[…]
+$\leq$ $C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\})(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}),$
 and
-$|\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}}.$
-Hence the proof of (3.4) is completed.
-Moreover, from (3.10),
-$\left[\partial_{\lambda_{R}}^{2}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
-$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
-$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}\partial^{2}_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
-- $\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
-$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}(4\pi\lambda_{I})\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
-${\color[rgb]{0,0,0}+}$ $\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi i\lambda_{I})\partial_{\lambda_{R}}\left(Gf\right).$
-Hence the proof of (3.5)-(3.7) can be justified by bootstrapping (3.10) or (3.11).
-Lemma 3.2.
-Suppose that (2.1) holds.
-(3.12) $|m_{0}-1|\leq C|u_{0}|_{\mathfrak{M}^{0,0}},\ \ |\partial_{\lambda_{R}}m_{0}|,\ |\partial_{\lambda_{I}}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|),$
-(3.13) $|\partial_{\lambda_{R}}^{j}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{j-1,0}}(1+|x_{2}|)^{j-1}(1+|\lambda_{I}|)^{j},\ j\leq 3,$
-(3.14) $|\partial_{\lambda_{I}}^{l}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{1-l,0}}(1+|x_{2}|)^{l-1}(1+|\lambda|)^{l},\ 2\leq l\leq 3,$
-(3.15) $|\partial_{\lambda_{R}}\partial_{\lambda_{I}}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{1,0}}(1+|x_{2}|)(1+|{\color[rgb]{0,0,0}\lambda}|)^{2}.$
-Proof.
-Estimate (3.12) follows from Lemma 3.1 and
-(3.16) $\begin{gathered}m_{0}=(1-G_{u_{0}})^{-1}1,\\ \partial_{\lambda_{R}}m_{0}=(1-G_{u_{0}})^{-1}\left[\partial_{\lambda_{R}}G_{u_{0}}\right]m_{0},\quad\partial_{\lambda_{I}}m_{0}=(1-G_{u_{0}})^{-1}\left[\partial_{\lambda_{I}}G_{u_{0}}\right]m_{0}.\end{gathered}$
-Applying the bootstrapping argument to (3.16) and Lemma 3.1, we can justify (3.13)-(3.15). ∎
-Combining Lemma 3.2 with (2.2), (2.10), and (2.12), we further deduce
-Lemma 3.3.
-Suppose the assumption of Theorem 1 holds for $\mathfrak{M}^{2,q}$.
-(3.17) $|\partial_{\lambda^{\prime}_{R}}^{j}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{j-1,0}}(1+|\lambda^{\prime}_{I}|)^{j},\ 1\leq j\leq 3,$
-(3.18) $|\partial_{\lambda^{\prime}_{I}}^{l}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{l-1,0}}(1+|\lambda^{\prime}|)^{l},\ \ 1\leq l\leq 3,$
-(3.19) $|\partial_{\lambda^{\prime}_{R}}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{1,0}}(1+|\lambda^{\prime}|)^{2}.$
+(3.26) $|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq C|\lambda_{I}\partial_{\lambda_{I}}s_{c}|$
+$\leq$ $C|\partial_{\lambda_{I}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{I}}\left[u_{0}(\partial_{x_{1}}m_{0})\right]^{\wedge}|$
+$\leq$ $C|\partial_{\lambda_{I}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{I}}\left[u_{0}(1-G_{u_{0}})^{-1}(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|$
+$\leq$ $C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\})(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}).$
 3.2. Long time asymptotics of $u_{1}(x)$
@@
 Proof.
-Integration by parts, applying Lemmas 3.3, 3.4, we have
+Integration by parts, applying Lemmas 3.2, 3.3, we have
 (3.42) $|u_{1,2}(x)|\leq$ $\frac{C}{t}|\iint e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3}-3\zeta^{\prime}_{I}{\zeta^{\prime}_{R}}^{2})}\nabla\cdot\left(\widetilde{s}_{c}(\zeta^{\prime})(\overline{\zeta}^{\prime}-\zeta^{\prime})(1-\chi)\frac{\nabla S_{0}}{|\nabla S_{0}|^{2}}\right)d\zeta^{\prime}_{R}d\zeta^{\prime}_{I}|,$
@@
 Proposition 3.2.
-Suppose the assumption of Theorem 1 holds for $\mathfrak{M}^{2,q}$. Then, as $t\to+\infty$:
+Suppose the assumption of Theorem 1 holds. Then, as $t\to+\infty$:
 (3.46) $u_{1,1}(x)\sim\frac{2ie^{i4\pi tr^{3}}}{3t}\widetilde{s}_{c}(+ir)-\frac{2ie^{-i4\pi tr^{3}}}{3t}\widetilde{s}_{c}(-ir)+\mathcal{O}(t^{-4/3}),\ \textit{for $\color[rgb]{0,0,0}a<-\delta<0$},$
@@
 (3.50) $u_{1,1}=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}e^{\pi i\frac{\operatorname{sgn}(\zeta^{\prime}_{I})}{4}}\frac{1}{\sqrt{|\frac{6}{\pi}\zeta^{\prime}_{I}|}}\int d\eta^{\prime}_{R}\left(1+\mathcal{O}(\frac{{\eta^{\prime}}^{2}_{R}}{t|\zeta^{\prime}_{I}|})\right)f^{\wedge_{\zeta^{\prime}_{R}}}(\eta^{\prime}_{R},\zeta^{\prime}_{I})$
-$=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}-{\zeta^{\prime}_{I}}^{3})}e^{+\pi i\frac{\operatorname{sgn}(\zeta^{\prime}_{I})}{4}}\frac{1}{\sqrt{|\frac{6}{\pi}\zeta^{\prime}_{I}|}}\psi_{r,r}(\zeta^{\prime}_{I})(\overline{\zeta}^{\prime}-\zeta^{\prime})\widetilde{s}_{c}(0,\zeta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}}).$
+$=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}e^{+\pi i\frac{\operatorname{sgn}(\zeta^{\prime}_{I})}{4}}\frac{1}{\sqrt{|\frac{6}{\pi}\zeta^{\prime}_{I}|}}\psi_{r,r}(\zeta^{\prime}_{I})(\overline{\zeta}^{\prime}-\zeta^{\prime})\widetilde{s}_{c}(0,\zeta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}}).$
 […]
@@
 Using (3.51), the Fourier multiplication formula, (3.53), and (3.55), (3.50) turns into
-(3.49) $u_{1,1}(x)=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\eta^{\prime}_{I}\left(e^{-2it(a\zeta^{\prime}_{I}-{\zeta^{\prime}_{I}}^{3})}\right)^{\wedge_{\zeta^{\prime}_{I}}}(-\eta^{\prime}_{I})g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}})$
+(3.56) $u_{1,1}(x)=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\eta^{\prime}_{I}\left(e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}\right)^{\wedge_{\zeta^{\prime}_{I}}}(-\eta^{\prime}_{I})g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}})$
 $=$ $-\frac{4}{(6t)^{\frac{1}{3}}}\frac{1}{\sqrt{t}}\int d\eta^{\prime}_{I}\ Ai\left(\frac{(2t)^{\frac{2}{3}}}{\sqrt[3]{3}}(a-\frac{\pi\eta^{\prime}_{I}}{t})\right)g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}}).$
@@
 Let $g_{+}=\psi_{r,0}(\zeta^{\prime}_{I})\partial_{\zeta^{\prime}_{R}}\left(\frac{1}{\zeta_{R}^{\prime}}\psi_{r,r}(\zeta^{\prime}_{R})\widetilde{s}_{c}(\zeta^{\prime})\right)$.
-Applying Lemma 3.3, and $u_{0}\in\mathfrak{M}^{1,q}$, we have
+Applying Lemma 3.1, and $\color[rgb]{0,0,0}u_{0}\in\mathfrak{M}^{2,q}$, we have
 $|g_{+}|_{L^{2}(d\zeta^{\prime}_{I})},\ |\partial_{\zeta^{\prime}_{I}}g_{+}|_{L^{2}(d\zeta^{\prime}_{I})},\ |g_{+}^{\wedge_{\zeta^{\prime}_{I}}}|_{L^{1}(d\eta^{\prime}_{I})}<C.$
@@
 Theorem 3.
-Suppose that (2.1) holds for $\mathfrak{M}^{2,q}$. Then, as $t\to+\infty$,
+Suppose that (2.1) holds. Then, as $t\to+\infty$,
 […]
@@
 (4.2) $\partial_{x_{1}}^{j}\widetilde{\mathcal{C}Tf}$= $-\frac{1}{2\pi i}\iint\frac{e^{2\pi itS_{0}}(2\pi i\xi_{1}^{\prime})^{j}\widetilde{s}_{c}(\zeta^{\prime})}{\zeta^{\prime}-\lambda^{\prime}}\widetilde{f}(\zeta^{\prime})d\overline{\zeta}^{\prime}\wedge d\zeta^{\prime}$
-${\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}\iint\frac{e^{2\pi itS_{0}(a;\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}(2\pi i\xi_{1}^{\prime})^{j}\widetilde{s}_{c}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi_{2}^{\prime})}\widetilde{f}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))d\xi^{\prime}_{1}d\xi^{\prime}_{2}},$
+${\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}-\iint\frac{e^{2\pi itS_{0}(a;\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}(2\pi i\xi_{1}^{\prime})^{j}{\color[rgb]{0,0,0}\left[u_{0}m_{0}\right]^{\wedge}(\xi_{1},\xi_{2},\zeta)}}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi_{2}^{\prime})}\widetilde{f}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))d\xi^{\prime}_{1}d\xi^{\prime}_{2}},$
 with
@@
 (4.6) $\mathfrak{S}(a;\lambda_{R}^{\prime};\xi_{1}^{\prime\prime})=4\pi^{2}\xi_{1}^{\prime\prime 3}+(a-3{\lambda^{\prime}}_{R}^{2})\xi_{1}^{\prime\prime},$
-$\mathcal{F}(t;\lambda^{\prime};x_{2}^{\prime};\xi_{1}^{\prime\prime})={\color[rgb]{0,0,0}\operatorname{sgn}(x_{2}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
+$\mathcal{F}(t;\lambda^{\prime};x_{2}^{\prime};\xi_{1}^{\prime\prime})={\color[rgb]{0,0,0}-\operatorname{sgn}(x_{2}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
 $\hskip 108.405pt\times e^{4\pi^{2}(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}.$
@@
 Using (2.12), Lemma 4.1, the Fourier transform theory, $\exp{(+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}}))}$ is holomorphic in $\xi^{\prime}_{2}$ when $\xi_{1}^{\prime}\neq 0$ (i.e., holomorphic in $\zeta^{\prime}_{R}\zeta_{I}^{\prime}$ when $\zeta_{I}^{\prime}\neq 0$), and the residue theorem, we formally derive
-(4.7) $\widetilde{\mathcal{C}T1}={\color[rgb]{0,0,0}\iint}[\,\frac{e^{+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})}}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi^{\prime}_{2})}\,]^{\vee_{\xi^{\prime}_{1},\xi^{\prime}_{2}}}(ta-x_{1}^{\prime},-x_{2}^{\prime})[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x_{2}^{\prime},x^{\prime}_{2})dx_{1}^{\prime}dx_{2}^{\prime},$
+(4.7) $\widetilde{\mathcal{C}T1}={\color[rgb]{0,0,0}-\iint}[\,\frac{e^{+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})}}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi^{\prime}_{2})}\,]^{\vee_{\xi^{\prime}_{1},\xi^{\prime}_{2}}}(ta-x_{1}^{\prime},-x_{2}^{\prime})[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x_{2}^{\prime},x^{\prime}_{2})dx_{1}^{\prime}dx_{2}^{\prime},$
 where $\mathfrak{m}_{0}$ satisfies (4.5) (see Lemma A.1 in the Appendix for the proof) and
@@
 $\frac{1}{2}\leq\beta_{n}=\frac{1}{2}(2-\beta_{n-1})\leq 1,\ \beta_{1}=1,$
-$\mathcal{F}^{(n)}(t;\lambda^{\prime};x_{2,n}^{\prime};\xi_{n}^{\prime\prime})={\color[rgb]{0,0,0}\operatorname{sgn}(x_{2,n}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
+$\mathcal{F}^{(n)}(t;\lambda^{\prime};x_{2,n}^{\prime};\xi_{n}^{\prime\prime})={\color[rgb]{0,0,0}-\operatorname{sgn}(x_{2,n}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
 $\times e^{\beta_{n}4\pi^{2}(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi})}.$
@@
 $\widetilde{(\mathcal{C}T)^{n}1}(t_{1},t_{2},t;\lambda^{\prime})$
-$=$ ${\color[rgb]{0,0,0}-\frac{1}{2\pi i}}\iint\left[\frac{e^{\frac{2-\beta_{n-1}}{2}2\pi iS_{0}(\zeta^{\prime})}\left[\mathfrak{C}\mathfrak{T}\right]^{(n-1)}(t_{1},t_{2},t;\overline{\zeta}^{\prime})}{p_{\lambda^{\prime}}({\xi^{\prime}_{1}},{\xi^{\prime}_{2}})}\right]^{\vee_{{\xi^{\prime}_{1}},{\xi^{\prime}_{2}}}}(ta-x_{1,n}^{\prime},-x_{2,n}^{\prime})$
+$=$ ${\color[rgb]{0,0,0}-}\iint\left[\frac{e^{\frac{2-\beta_{n-1}}{2}2\pi iS_{0}(\zeta^{\prime})}\left[\mathfrak{C}\mathfrak{T}\right]^{(n-1)}(t_{1},t_{2},t;\overline{\zeta}^{\prime})}{p_{\lambda^{\prime}}({\xi^{\prime}_{1}},{\xi^{\prime}_{2}})}\right]^{\vee_{{\xi^{\prime}_{1}},{\xi^{\prime}_{2}}}}(ta-x_{1,n}^{\prime},-x_{2,n}^{\prime})$
 $\times$ $[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1,n}-\frac{2t_{2}}{3}x^{\prime}_{2,n},x^{\prime}_{2,n})dx_{1,n}^{\prime}dx_{2,n}^{\prime}$
@@
 Lemma 4.6.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (4.48) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}[1-\chi(\lambda^{\prime})]}$
@@
 Lemma 4.7.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (4.54) $\hskip 36.135pt\sum_{n=1}^{\infty}\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\chi(\lambda^{\prime})}\\ \times\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}\sim{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.\hskip 36.135pt$
@@
 Theorem 4.
-Assume the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,
+Assume the assumption of Theorem 1 holds. As $t\to+\infty$,
 (4.55) $u_{2,0}\sim{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.$
@@
 Lemma 4.8.
-Suppose the assumption of Theorem 1 is valid for $u_{0}\in\mathfrak{M}^{3,q}$. Then there exists a constant $C_{0}\gg 1$ such that, defining
+Suppose the assumption of Theorem 1 is valid. Then there exists a constant $C_{0}\gg 1$ such that, defining
 (4.56) $\color[rgb]{0,0,0}\Xi(\lambda^{\prime})=\theta(C_{0}-|\lambda^{\prime}_{R}|)+\theta(|\lambda^{\prime}_{R}|-C_{0})\theta(|\lambda^{\prime}_{I}|-1),$}
@@
 Lemma 4.9.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (4.67) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\Xi(\lambda^{\prime})\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
@@
 (4.68) ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|{\lambda^{\prime}}_{R}|-t^{-5/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})$
-$=$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|{\lambda^{\prime}}_{R}|-t^{-5/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-3.5/9})$
-$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-1.9/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
-$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)$
-$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
-$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|).$
-From $\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})\neq 0$}, we can show that the $L^{1}(d\xi^{\prime\prime}_{n})$-norms of the amplitude function $\mathcal{F}^{(n)}$ over the corresponding domains for the first and second terms on the right-hand side of (4.68) are $o(t^{-1})$.
-Along with (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, and (2.12), the analysis reduces to studying the contribution over the domain corresponding to the third term, which is bounded by:
+$=$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-1.9/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})$
+$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9}).$
+From $\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})\neq 0$}, we can show that the $L^{1}(d\xi^{\prime\prime}_{n})$-norms of the amplitude function $\mathcal{F}^{(n)}$ over the corresponding domains for the first term on the right-hand side of (4.68) is $o(t^{-1})$.
+Along with (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, and (2.12), the analysis reduces to studying the contribution over the domain corresponding to the second term, which is bounded by:
 (4.69) $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
-$\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
-$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
+$\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
 $+$ $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
 $\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)(1-\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime}))\mathfrak{C}\mathfrak{T}_{0,(n)}\psi_{t^{-0.95/3},b}(\xi_{n}^{\prime\prime})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
-$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
+$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
 $+$ $|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
 $\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)(1-\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime}))\mathfrak{C}\mathfrak{T}_{0,(n)}(1-\psi_{t^{-0.95/3},b}(\xi_{n}^{\prime\prime}))\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
-$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
-$\equiv$ $I_{1}+I_{2}+I_{3}.$
+$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|\equiv I_{1}+I_{2}+I_{3}.$
 Applying Proposition 4.1, and using estimates $|\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime})|_{L^{1}(d\lambda_{I}^{\prime})}\leq Ct^{-0.9/3}$, $|\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)|_{L^{1}(d\lambda_{R}^{\prime})}\leq Ct^{-1.9/9}$, and (2.3), we obtain
@@
 Lemma 4.10.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (4.75) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\widetilde{s}_{c}(\lambda^{\prime})e^{\beta_{n+1}2\pi itS_{0}}(\overline{\lambda}^{\prime}-\lambda^{\prime}){\color[rgb]{0,0,0}\Xi(\lambda^{\prime})\psi_{r/C,0}(\lambda^{\prime}_{R})\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
@@
 If the assumption of Theorem 1 holds for $u_{0}\in{\mathfrak{M}^{1,q}}$ then
-(5.1) $\partial_{x_{1}}\widetilde{\mathcal{C}T1}(t_{1},t_{2},t;\lambda^{\prime})$
-$=$ $e^{i\pi tS_{0}(a;\lambda^{\prime})}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ \left(\partial_{x^{\prime}_{1}}[u_{0}\mathfrak{m}_{0}]\right)(x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
-$\times$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{1}^{\prime},x_{2}^{\prime};\xi_{1}^{\prime\prime})\equiv e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{1,(1)}1,$
+(5.1) $\partial_{x_{1}}\widetilde{\mathcal{C}T1}(t_{1},t_{2},t;\lambda^{\prime})=e^{i\pi tS_{0}(a;\lambda^{\prime})}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ \left(\partial_{x^{\prime}_{1}}[u_{0}\mathfrak{m}_{0}]\right)(x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
+$\quad\times\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{1}^{\prime},x_{2}^{\prime};\xi_{1}^{\prime\prime})\equiv e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{1,(1)}1,$
 or
@@
 is holomorphic in $\lambda^{\prime}_{R}\lambda^{\prime}_{I}$ when ${\lambda^{\prime}}_{I}\neq 0$. Here
-(5.4) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
-$=$ $\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\mathfrak{C}\mathfrak{T}_{1,(h)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
+(5.4) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})=\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\mathfrak{C}\mathfrak{T}_{1,(h)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
 or
-(5.5) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
-$=$ $\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}$
-$\times$ $\{\,\mathfrak{C}\mathfrak{T}_{1,(h)}[1-\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})]+\mathfrak{C}\mathfrak{T}_{0,(h)}\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})\cdot{\color[rgb]{0,0,0}(2\pi i)}(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\,\}$
-$\times$ $\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
+(5.5) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})=\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}$
+$\quad\times\{\,\mathfrak{C}\mathfrak{T}_{1,(h)}[1-\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})]+\mathfrak{C}\mathfrak{T}_{0,(h)}\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})\cdot{\color[rgb]{0,0,0}(2\pi i)}(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\,\}$
+$\quad\times\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
 where $\xi^{\prime\prime}_{n+1}=\frac{\lambda_{I}^{\prime}}{2\pi}$.
@@
 Proposition 5.2.
-Suppose the assumption of Theorem 1 is valid for $u_{0}\in\mathfrak{M}^{3,q}$, $\Xi(\lambda^{\prime})$ defined by (4.56), and $|a|>\delta>0$. Then, as $t\to\infty$,
+Suppose the assumption of Theorem 1 is valid, $\Xi(\lambda^{\prime})$ defined by (4.56), and $|a|>\delta>0$. Then, as $t\to\infty$,
 (5.8) $|u_{2,1}(x)|\leq$ $C\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
@@
 Lemma 5.2.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (5.10) $u_{2,1}(x)\leq$ $C\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
@@
 Lemma 5.3.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (5.16) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9}){\color[rgb]{0,0,0}\theta(a-3{\lambda^{\prime}_{R}}^{2})}$
@@
 Lemma 5.4.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (5.17) $\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(a-3{\lambda^{\prime}_{R}}^{2})}$
@@
 Lemma 5.5.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (5.25) $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime}){\color[rgb]{0,0,0}\theta(a-3{\lambda^{\prime}_{R}}^{2})}$
@@
 Theorem 6.
-Assume the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,
+Assume the assumption of Theorem 1 holds. As $t\to+\infty$,
 (5.29) $u_{2,1}\sim{\color[rgb]{0,0,0}\mathcal{O}(t^{-1})}.$
@@
 Lemma 5.6.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 (5.30) $u_{2,1}(x)\leq$ $C\sum_{n=1}^{\infty}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})$
@@
 Lemma 5.7.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathbb{P}_{n,h}^{>}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-8/9}).$
@@
 Lemma 5.8.
-Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
+Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
 $\sum_{n=1}^{\infty}\sum_{h=1}^{n}|\iint d\overline{\lambda}^{\prime}\wedge d\lambda^{\prime}\ e^{\beta_{n+1}2\pi itS_{0}}\widetilde{s}_{c}(\lambda^{\prime})\Xi(\lambda^{\prime})\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\mathbb{P}_{n,h}^{<}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}|\leq o(t^{-8/9}).$
@@
 Theorem 7.
-Assume the assumption of Theorem 1 holds for $u\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,
+Assume the assumption of Theorem 1 holds. As $t\to+\infty$,
 $u_{2,1}\sim o(t^{-8/9}).$